### v1

# Planning behavior in a recurrent neural network that plays Sokoban

Abstract
To predict how advanced neural networks generalize to novel situations, it is essential to understand how they reason. Guez et al. (2019, “An investigation of model-free planning”) trained a recurrent neural network (RNN) to play Sokoban with model-free reinforcement learning. They found that adding extra computation steps to the start of episodes at test time improves the RNN’s success rate. We further investigate this phenomenon, finding that it rapidly emerges early on in training and then slowly fades, but only for comparatively easier levels. The RNN also often takes redundant actions at episode starts, and these are reduced by adding extra computation steps. Our results suggest that the RNN learns to take time to think by ‘pacing’, despite the per-step penalties, indicating that training incentivizes planning capabilities. The small size (1.29M parameters) and interesting behavior of this model make it an excellent model organism for mechanistic interpretability.
Machine Learning, ICML
Figure 1: Top: Proportion of medium-difficulty validation levels solved vs. environment steps used in training. Curves show the RNN’s performance with a specific number of forced thinking steps at episode start, along with a ResNet baseline. Bottom: Estimated planning effect: the 8-steps minus the 0-steps curve. We see that reasoning emerges in the first 70M steps and keeps increasing for the hard levels, but decreases for medium levels.
1 Introduction
In many tasks, performance of both humans and sufficiently advanced neural networks (NNs) improves with more reasoning: whether giving a human more time to think before taking a chess move, or prompting a large language model to reason step by step (Kojima et al., 2022). Thus, understanding reasoning in such NNs is key to predicting their behavior in novel situations.
Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment. So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning (Hubinger et al., 2019) – may internalize goals different from that given by the training objective, leading to harmful misgeneralization (Di Langosco et al., 2022; Shah et al., 2022).
This paper presents a 1.29M parameter recurrent neural network (RNN) that clearly benefits from extra time to reason. Following Guez et al. (2019), we train a convolutional LSTM to play Sokoban, a challenging puzzle game that remains a benchmark for planning algorithms (Peters et al., 2023) and reinforcement learning (Chung et al., 2024).
We believe understanding this RNN is a useful first step towards understanding reasoning in sophisticated neural networks. The network is complex enough to be interesting: it reasons in a challenging environment and benefits from extra thinking time. At the same time, it is simple enough to be tractable to reverse engineer with current mechanistic interpretability techniques, owing to its small size and relatively straightforward behavior.
We make three key contributions. First, we replicate and open-source an RNN whose performance improves with thinking time (Guez et al., 2019). Second, we analyze its training dynamics, finding that the thinking-time behavior emerges early in training and is incentivized by the setup (Figure 1, Sections 3 and 4.1). Third, we conduct a detailed empirical analysis of the network’s behavior. We quantitatively investigate how increased thinking changes network behavior (Section 3), showing that it reduces myopic actions (Figure 5). Furthermore, we present case studies of planning behavior that suggest the RNN naturally takes time to think (Section 4), which may explain the reduced planning effect on medium levels (Figure 1).
2 Training the test subject
We closely follow the setup from Guez et al. (2019), using the IMPALA (Espeholt et al., 2018) reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture, as well as a ResNet. We open-source both the trained networks [^1] https://github.com/AlignmentResearch/learned-planner and training code [^2] https://github.com/AlignmentResearch/train-learned-planner.
Dataset.
Sokoban is a grid puzzle game with walls, floors, movable boxes, and target tiles. The player’s goal is to push all boxes onto target tiles while navigating walls. We use the Boxoban dataset (Guez et al., 2018), consisting of $10\times 10$ procedurally generated Sokoban levels, each with 4 boxes and targets. The edge tiles are always walls, so the playable area is $8\times 8$. Boxoban separates levels into train, validation and test sets, with three difficulty levels: unfiltered, medium, and hard. Guez et al. (2019) generated these by filtering levels that cannot be solved by progressively better-trained DRC networks, so easier sets occasionally contain difficult levels. In this paper, we use unfiltered-train (900k levels) to train networks. To evaluate them, we use unfiltered-test (1k levels) [^3] We prefer unfiltered-test rather than unfiltered-validation so the numbers are directly comparable to Guez et al. (2019)., medium-validation (50k levels), and hard ($\sim$3.4k levels), which do not overlap.
Environment.
The observations are $10\times 10$ RGB images, normalized by dividing each component by 255. Each type of tile is represented by a pixel of a different color (Schrader, 2018). See Figure 6 for examples. The player can move in cardinal directions (Up, Down, Left, Right). The reward is -0.1 per step, 1 for putting a box on a target, -1 for removing a box, and 10 for finishing the level. To avoid strong time correlations during learning, each episode resets at a uniformly random length between 91 and 120 time steps.
Figure 2: Left: A Sokoban level from the hard set. Right: the same level as the NN sees it, one pixel per tile. Walls are black, boxes are brown, targets are pink and the robot is green.
$\text{DRC}(D,N)$ architecture.
Guez et al. (2019) introduced the Deep Repeating ConvLSTM (DRC), whose core consists of $D$ convolutional LSTM layers with 32 channels and $3\times 3$ filters, each applied $N$ times per time step. Our $\text{DRC}(3,3)$ – or just DRC for brevity – has $1.29$M parameters. Before the LSTM core, two convolutional layers (without nonlinearity) encode the observation with $4\times 4$ filters.
The LSTM core uses $3\times 3$ convolutional filters, and a nonstandard tanh on the output gate (Jozefowicz et al., 2015). Unlike the original ConvLSTM (Shi et al., 2015), the input to each layer of a DRC consists of several concatenated components:
  • The encoded observation is fed into each layer.
  • To allow spatial information to travel fast in the ConvLSTM layers, we apply pool-and-inject by max- and mean-pooling the previous step’s hidden state. We linearly combine these values channel-wise before feeding them as an input to the next step.
  • To avoid convolution edge effects from disrupting the LSTM dynamics, we feed in a $12\times 12$ channel with zeros on the inside and ones on the boundary. Unlike the other inputs, this one is not zero-padded, maintaining the output size.
ResNet architecture.
This is a convolutional residual neural network, also from Guez et al. (2019). It serves as a non-recurrent baseline that can only think during the forward pass (no ability to think for extra steps) but is nevertheless good at the game. The ResNet consists of 9 blocks, each with $4\times 4$ convolutional filters. The first two blocks have 32 channels, and the others have 64. Each block consists of a convolution, then two (relu, conv) sub-blocks, which each split off and are added back to the trunk. The ResNet has $3.07$M parameters.
Value and policy heads.
After the convolutions, an affine layer projects the flattened spatial output into 256 hidden units. We then apply a ReLU and two different affine layers: one for the actor (policy) and one for the critic (value function).
RL training.
[…]
A* solver.
We used the A* search algorithm to obtain optimal solutions to each Sokoban puzzle. The heuristic was the sum of the Manhattan distances of each box to its nearest target. Solving a single level on one CPU takes anywhere from a few seconds to 15 minutes. [^5] The A* solutions may be of independent interest, so we make them available at https://huggingface.co/datasets/AlignmentResearch/boxoban-astar-solutions/.
3 Quantitative behavior analysis
Figure 3: Proportion of medium-difficulty validation levels solved by DRC with $x$ time steps to think. The dotted line shows the ResNet’s solve proportion. The proportion of solved levels increases with thinking time, but only up to 6 steps.
3.1 Recurrent networks are better at Sokoban
After training both the recurrent $\text{DRC}(3,3)$ network and feed-forward ResNet for 2B environment transitions, the $\text{DRC}(3,3)$ network is able to solve $99.3\%$ of unfiltered test levels, while the ResNet solves only $97.9\%$. This gap is particularly notable given that the ResNet has more than double the parameters of the $\text{DRC}(3,3)$ network, so the ResNet has more model capacity and sequential thinking time. These results are consistent with Guez et al. (2019), see Appendix B.
The performance gap widens when considering just the hard levels, of which the $\text{DRC}(3,3)$ network solves $42.8\%$ and the ResNet $26.2\%$. Counterintuitively, we find that – controlling for performance on the unfiltered test set – recurrent networks are not comparatively better at harder levels than the ResNet. Figure 10 plots the performance of model checkpoints on the unfiltered test set against medium-difficulty and hard levels. We find that both model architectures follow the same curve: the recurrent $\text{DRC}(3,3)$ network is further to the right, but follows the same trend.
3.2 Thinking time improves performance
We find the performance of our trained $\text{DRC}(3,3)$ improves given additional thinking time, replicating Guez et al. (2019). We evaluate the DRC network on all medium-difficulty validation levels. To induce thinking time, we repeat the initial environment observation $n$ times while advancing the DRC hidden state, then let the DRC policy act normally. Figure 3 shows the DRC’s performance for $n\in\{0,1,2,4,6,8,10,12,16\}$ steps, alongside the ResNet baseline. The DRC’s success rate improves from $76.6\%$ to $81.3\%$ with more thinking steps, with peak performance at 6 steps. The effect ($4.7\%$ difference) is comparable to that observed by Guez et al. ($4.5\%$).
Figure 4: Average optimal solution length of levels grouped by the number of thinking steps at which the level is first solved. Each of the 50,000 medium-difficulty validation levels is assigned to a group. Optimal solution length is a proxy for difficulty, and we can see that levels which take longer to solve tend to be harder.
3.3 Planning emerges early in training
Figure 1 shows the performance of the DRC on the medium-difficulty validation levels by training step. We see the DRC rapidly acquires a strong planning effect by 70M environment steps. This effect strengthens on the hard levels as training progresses, but slowly weakens for medium-difficulty levels. This suggests that the training setup and loss function reinforce the planning effect for harder levels and disincentivize it for levels that are comparatively easier.
3.4 Do difficult levels benefit more from thinking?
The A* solver gives us two proxies for the difficulty of a level: the length of the optimal solution, and the number of nodes it had to expand to find a solution. Figure 4 shows a positive correlation between the A* optimal solution length ($y$-axis) and the number of DRC thinking steps $n$ needed to solve a level ($x$-axis). In other words, taking extra more thinking time helps solve harder levels.
On the other hand, Figure 11 (Appendix D) finds only a weakly positive correlation between the expanded nodes proxy and required thinking steps $n$, with results noisy for smaller $n$. We speculate that the ‘expanded nodes’ proxy metric is only weakly related to required $n$ due to the DRC having different heuristics than the one we use for A*, whereas the optimal solution length does not depend on heuristics.
Table 1: Levels at 6 thinking steps.
Level categorization Percentage
Solved, previously unsolved 6.87
Unsolved, previously solved 2.23
Solved, with better returns 18.98
Solved, with the same returns 50.16
Solved, with worse returns 5.26
Unsolved, with same or better returns 15.14
Unsolved, with worse returns 1.36
3.5 Other consequences of DRC thinking time
While Figure 3 shows that the DRC can solve many more levels when given thinking steps at the beginning of an episode, in this section, we explore how the network’s behavior changes with thinking time, and if this has significant effects besides solving more levels.
Thinking makes the DRC “patient.”
One hypothesis is that, without the extra thinking steps, the DRC sometimes comes up with a greedy solution for some of the boxes and executes it right away. This puts some boxes into the wrong targets, and makes the puzzle unsolvable. Since the training signal penalizes every step it takes to solve the puzzle, it could be optimal for a bounded rational agent to avoid thinking and be “impatient” in this way: the agent may accumulate more return on average by being “impatient” and quickly solving most levels even if it sometimes errs by taking actions that make it impossible to solve other levels.
For $n$ thinking steps, Figure 5 plots the average number of steps (time-to-box) until each of 4 boxes gets on a target, in the order in which they do so. The results support the hypothesis: levels that are solved at 6 thinking steps but unsolved at 0 thinking steps show increased time-to-box for boxes 1-3 but decreased time for box 4 (Figure 5(b)). In contrast, averaged across all levels times go down for all boxes with increased thinking (Figure 5(a)).
Thinking often improves solutions for already-solved levels, but is sometimes harmful.
Table 1 shows that an additional 6 thinking steps solves 6.87% of previously unsolved levels – but at the cost of no longer being able to solve 2.23% of previously solved levels, for a total improved solution rate of 4.64%. In addition to improving the solution rate, additional thinking improves returns on average by shortening solution length in 18.98% of episodes, consistent with the decreased timesteps to place boxes shown in Figure 5(a). However, the effect is inconsistent: most episodes show no change in returns, with returns even decreasing in 5.26% of episodes. Overall, extra thinking steps are not always beneficial but, on balance, they let the DRC solve more levels or improve return on already solved levels.
Figure 5: Average time step to place each box $Bi$ on target for different thinking steps. (a) Averages across all levels where the box $Bi$ is placed on target by DRC with $n$ thinking steps. If box $Bi$ is not placed (on unsolved levels) we ignore the box, but we keep the other boxes in the level. (b) Averages for all levels solved by $6$ thinking steps but not solved by 00 thinking steps.
(a) file 0, level 18
[…]
(b) file 0, level 53
[…]
(c) file 0, level 153
[…]
Figure 6: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps. Colors are as in Figure 2. Boxes and targets are paired in upper- and lower-case letters respectively, such that the DRC’s best solution places boxes on targets in alphabetical order. Videos available at this https URL; see Section 4 for behavior descriptions. Levels solved faster obtain higher return as the agent incurs the per-step penalty fewer times.
4 Case studies: how does thinking change behavior?
To gain a deeper understanding of the DRC networks’ behavior, we examine an example from each of three categories from Table 1: a solved, previously unsolved level; a level on which the DRC improved its return with extra thinking (solved faster); and a level on which the return worsens (solved slower). We select an illustrative example of each category from the first ten examples by level index.
Thinking lets the DRC solve: Figure 6(a).
In the no-thinking condition, the DRC first pushes box $C$ one square to the right. It then goes back to push $A$ to $a$, but it is too late: it is now impossible to push box $B$ onto $b$. The network pushes $C$ onto $c$, then $D$ onto $d$ and then remains mostly still. In contrast, after thinking, the DRC pushes $A$ to $a$ first, which lets it solve the level.
Thinking speeds up solving: Figure 6(b).
In the no-thinking condition, the DRC spends many steps going back and forth before pushing any boxes. First it goes down to $y$, then up to $c$, then down onto $z$, back up to $y$ and to $z$ again. It then proceeds to solve the rest of the puzzle correctly: push box $A$ onto $a$, prepare box $B$ on $x$ and box $C$ where $B$ originally was, push in boxes $B,C$ and finally $D$. In the thinking condition, the DRC makes a beeline for $A$ and then plays the exact same solution.
Thinking slows down solving: Figure 6(c).
In the no-thinking condition, the DRC starts by pushing box $C$ into position $y$, then pushes boxes $A,B$. On the way back down, ther DRC pushes $C$ onto $c$ and finally $D$ onto $d$. In contrast, after thinking for a bit, the DRC goes the other way and starts by pushing $B$ onto $b$. The solution is the same after that: $A$, $C$, then $D$; but the NN has wasted time trekking back from $B$ to $A$.
4.1 Hypothesis: the DRC ‘paces’ to get thinking time
The second case study above found the DRC sometimes moves around in cycles without touching any box. One hypothesis for why this occurs is that the DRC is using the extra steps to come up with a plan to solve the episode. To test this, we can check whether cycles in the game state occur more frequently near the start of the episode, and whether deliberately giving the DRC thinking time makes it stop going in cycles.
Consistent with this hypothesis, Figure 7 (left) shows the majority of cycles start within the first 5 steps of the episode, and (right) forcing it to think for six steps makes about 75% of these cycles disappear.
The episode start is not the only time at which the DRC moves in cycles. Figure 9 shows that replacing an $n$-length cycle with $n$ thinking steps leads the NN to have the exact same behavior for at least 60% of levels, for at least 30 steps after the cycle. For context, the median solution length for train-unfiltered is exactly 30 steps. Additionally, in $82.39\%$ of cases, doing this prevents cycles from starting in the $n$ steps after the thinking-steps conclude (Appendix C).
Why does this behavior emerge during training? Thinking is useful for achieving higher return, so it should be reinforced. But it also has a cost, -0.1 per step, so it should be discouraged in easy levels that do not need the computation. We speculate that, as training advances and heuristics get tuned, the DRC needs to think in fewer levels, and it is better at knowing when it needs to pace. This would also explain why, after many training steps, the DRC network begins to benefit less from additional thinking steps at the start of the episode in Figure 1: the DRC finds levels easier and also knows to pace when thinking is needed.
Figure 7: Left: Histogram of cycle start times on the medium-difficulty validation levels. Right: Total number of cycles the agent takes in the first 5 steps across all episodes in medium-difficulty validation levels with $n$ initial thinking steps.
5 Related work
Interpretability of agents and world models.
Several works have attempted to find the mechanism by which a simple neural network does planning in mazes (Ivanitskiy et al., 2023; Mini et al., 2023), gridworlds (Bloom & Colognese, 2023), and graph search (Ivanitskiy et al., 2023). We believe the DRC we present is a clearer example of an agent than what these works focus on, and should be similarly possible to interpret.
Other works have found emergent world models in sequence-prediction (Li et al., 2023) and navigation (Wijmans et al., 2023) neural networks.
Goal misgeneralization and mesa-optimizers for alignment.
From the alignment perspective, AIs optimizing monomaniacally for a goal have long been a concern (e.g. Yudkowsky, 2006; Omohundro, 2008; see the preface of Russell, 2019). In a machine learning paradigm (Hubinger et al., 2019), the goal of the training system is not necessarily optimized; instead, the NN may optimize for a related or different goal (Di Langosco et al., 2022; Shah et al., 2022), or for no goal at all.
Chain-of-thought faithfulness.
Large language models (LLMs) use chain of thought, but are they faithful to it or do they think about their future actions in other ways (Lanham et al., 2023; Pfau et al., 2024)? One could hope that LLMs perform all long-term reasoning in plain English, allowing unintended human consequences to be easily monitored, as in Scheurer et al. (2023).
Reasoning neural network architectures.
Many papers try to enhance NN thinking by altering the training setup or architecture (Bansal et al., 2022; Graves, 2016; Chung et al., 2024).
Ethical treatment of AIs.
Do AIs deserve moral consideration? Schwitzgebel & Garza (2015) argue that very human-like AIs are conceivable and clearly deserve rights. Tomasik (2015) suggests that most AIs deserve at least a little consideration, like biological organisms of any species (Singer, 2004). But what does it mean to treat an AI ethically? Daswani & Leike (2015) argue that the way to measure pleasure and pain in a reinforcement learner is not by its absolute amount of return, but rather by the temporal difference (TD) error: the difference between its expectations and the actual return it obtained. If the internals of the NN have a potentially different objective (Hubinger et al., 2019; Di Langosco et al., 2022), then the TD error should come from a place other than the critic. This paper is an early step toward finding the learned-reward internal TD error, if it exists.
6 Conclusion
We have developed and open-sourced a model organism for understanding reasoning in neural networks. In addition, we contribute a detailed empirical analysis of this network’s behavior. We believe this work will prove useful to the interpretability, alignment, and ethical treatment of AIs.
First, we have trained a DRC recurrent network that plays Sokoban and benefits from additional thinking steps at test time, replicating Guez et al. (2019).
Second, we have shown that thinking for longer helps solve harder levels, and makes the DRC better at levels that require longer-sighted behavior. Without intervention, thinking sometimes takes the form of the DRC ‘pacing’ at the beginning or middle of the episode, in a way which can be substituted by repeating the same input. This suggests the network is deliberately using more computation by pacing.
Finally, we have shown that the training setup incentivizes the planning effect at the start. We find that planning is disincentivized later, but only for easier levels. We offer a hypothesis about why the training process disincentivizes the planning effect: the NN finds levels easier (needs less thinking), and also learns when to do the thinking it needs (via pacing).
7 Acknowledgements
We would like to thank ChengCheng Tan for editing this paper, Philip Quirke for project management, and the rest of the FAR team for general support. We also thank Maximilian Li, Alex F. Spies, and anonymous reviewers for improving the paper with their comments and edits. We are also very grateful for the many questions about their RL setup that Arthur Guez and Timothy Lillicrap answered, as well as inspiration from Stephen Chung and Scott Goodfriend.
Appendix A Training hyperparameters
All networks were trained with the same hyperparameters, which were tuned on a combination of the ResNet and the $\text{DRC}(3,3)$. These are almost exactly the same as Guez et al. (2019), allowing for taking the mean of the per-step loss instead of the sum.
Loss.
The value and entropy coefficients are 0.25 and 0.01 respectively. It is very important to not normalize the advantages for the policy gradient step.
Gradient clipping and epsilon
The original IMPALA implementation, as well as Huang et al. (2023), sum the per-step losses. We instead average them for more predictability across batch sizes, so we had to scale down some parameters by a factor of $1/640$: Adam $\epsilon$, gradient norm for clipping, and L2 regularization).
Weight initialization.
We initialize the network with the Flax (Heek et al., 2023) default: normal weights truncated at 2 standard deviations and scaled to have standard deviation $\sqrt{1/\text{fan\_in}}$. Biases are initialized to 0. The forget gate of LSTMs has 1 added to it (Jozefowicz et al., 2015). We initialize the value and policy head weights with orthogonal vectors of norm 1. Surprisingly, this makes the variance of these unnormalized residual networks decently close to 1.
Adam optimizer.
As our batch size is medium-sized, we pick $\beta_{1}=0.9$, $\beta_{2}=0.99$. The denominator epsilon is $\epsilon=1.5625\cdot 10^{-7}$. Learning rate anneals from $4\cdot 10^{-4}$ at the beginning to $4\cdot 10^{-6}$ at 2,002,944,000 steps.
L2 regularization.
In the training loss, we regularize the policy logits with L2 regularization with coefficient $1.5625\times 10^{-6}$. We regularize the actor and critic heads’ weights with L2 at coefficient $1.5625\times 10^{-8}$. We believe this has essentially no effect, but we left it in to more closely match Guez et al. (2019).
Software.
We base our IMPALA implementation on Cleanba (Huang et al., 2023). We implemented Sokoban in C++ using Envpool (Weng et al., 2022) for faster training, based on gym-sokoban (Schrader, 2018).
A.1 Number of training steps
[…]
Appendix B Learning curve comparison
It is difficult to fully replicate the results by Guez et al. (2019). Chung et al. (2024) propose an improved method for RL in planning-heavy domains. They employ the IMPALA $\text{DRC}(3,3)$ as a baseline and plot its performance in Chung et al. (2024, Figure 5). They plot two separate curves for $\text{DRC}(3,3)$: that from Guez et al. (2019), and a decent replicated baseline. The baseline is considerably slower to learn and peaks at lower performance.
We did not innovate in RL, so were able to spend more time on the replication. We compare our replication to Guez et al. (2019) in Figure 8, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1). Our implementation also appears much less stable, with large error bars and large oscillations over time. We leave addressing that to future work.
The success rate in Figure 8 is computed over 1024 random levels, unlike the main body of the paper. Table 3 reports test and validation performance for the DRC and ResNet seeds which we picked for the paper body.
The parameter counts (Table 2) are very different from what Guez et al. (2019) report. In private communication with the authors, we confirmed that our architecture has a comparable number of parameters, and some of the originally reported numbers are a typographical error.
Table 2: Parameter counts for each architecture.
Architecture Parameter count
$\text{DRC}(3,3)$ 1,285,125 (1.29M)
DRC(1,1) 987,525 (0.99M)
ResNet: 3,068,421 (3.07M)
Figure 8: Success rate for Test-unfiltered and Validation-medium levels vs. environment steps of training. Each architecture has 5 random seeds, the solid line is the pointwise median and the shaded area spans from the minimum to the maximum. The dotted lines are data for the performance of architectures extracted from the (Guez et al., 2019) PDF file. The values are slightly different from what Figures 1 and 3 report because they are calculated on a random sample of 1024 levels (24 levels are repeated for test-unfiltered).
Table 3: Success rate and return of DRC and ResNet on the unfiltered test set at various training environment steps.
Training Env Test Unfiltered Valid Medium
Steps ResNet DRC(3, 3) ResNet DRC(3, 3)
Success Return Success Return Success Return Success Return
100M 87.8 8.13 95.4 9.58 18.6 -6.59 47.9 -0.98
500M 93.1 9.24 97.9 10.21 39.7 -2.64 66.6 2.62
1B 95.4 9.75 99.2 10.47 50.0 -0.64 70.4 3.40
2B 97.9 10.29 99.3 10.52 59.4 1.16 76.6 4.52
Appendix C Additional cycles investigation
We run the DRC(3,3) on medium-validation levels and record where cycles in the game state happen. We prune out the redundant cycles that visit the same states as a given cycle. There are in total 13702 non-redundant cycles. For each cycle, if it is of length $N$, we replace it with $N$ steps of thinking, and measure what happens. In 86.74% of cases the DRC(3,3) does not immediately start going in cycles after the thinking time. That is, no cycle begins in the step after thinking time. In 82.39% of cases this persists for the next $N$ time steps: no cycles start in the $N$ steps after thinking.
(a) Cycle length distribution
(b) with cycles from all levels
(c) with cycles from those levels that were solved by the agent
Figure 9: We replace N-length cycles with N thinking steps and checking for the same state after some timesteps. (a) A histogram of cycle lengths in the medium-validation set. (b, c) After replacing a cycle with the same length in thinking steps, are all the states the same for the next $x$ steps?
Appendix D Additional quantitative behavior figures
Figure 10: Success rate on datasets of various difficulty, for various checkpoints of each architecture. This deviates very little from a curve, which shows that ResNets and DRCs which are equally good at the easier sets are also equally good at the harder sets. Perhaps DRC(1,1) is a slight exception, but it also performs much worse than the others overall (see Figure 8).
Figure 11: Number of thinking steps required to solve the level vs. number of nodes A* needs to expand to solve it. The trend is somewhat increasing but much less clear, indicating different heuristics used by the NN and A*.

### v2

@@
-Planning behavior in a recurrent neural network that plays Sokoban
+Planning in a recurrent neural network that plays Sokoban
 Abstract
-To predict how advanced neural networks generalize to novel situations, it is essential to understand how they reason.
+How a neural network (NN) generalizes to novel situations depends on whether it has learned to select actions heuristically or via a planning process.
 Guez et al.
-(2019, “An investigation of model-free planning”) trained a recurrent neural network (RNN) to play Sokoban with model-free reinforcement learning.
-They found that adding extra computation steps to the start of episodes at test time improves the RNN’s success rate.
-We further investigate this phenomenon, finding that it rapidly emerges early on in training and then slowly fades, but only for comparatively easier levels.
-The RNN also often takes redundant actions at episode starts, and these are reduced by adding extra computation steps.
-Our results suggest that the RNN learns to take time to think by ‘pacing’, despite the per-step penalties, indicating that training incentivizes planning capabilities.
-The small size (1.29M parameters) and interesting behavior of this model make it an excellent model organism for mechanistic interpretability.
-Machine Learning, ICML
-Figure 1: Top: Proportion of medium-difficulty validation levels solved vs.
+[2019, “An investigation of model-free planning”] found that a recurrent NN (RNN) trained to play Sokoban appears to plan, with extra computation steps improving the RNN’s success rate.
+We replicate and expand on their behavioral analysis, finding the RNN learns to give itself extra computation steps in complex situations by “pacing” in cycles.
+Moreover, we train linear probes that predict the future actions taken by the network and find that intervening on the hidden state using these probes controls the agent’s subsequent actions.
+Leveraging these insights, we perform model surgery, enabling the convolutional NN to generalize beyond its $10\times 10$ architectural limit to arbitrarily sized inputs.
+The resulting model solves challenging, highly off-distribution levels.
+We open-source our model and code, and believe the neural network’s small size (1.29M parameters) makes it an excellent model organism to deepen our understanding of learned planning.
+[…]
+Figure 1: Top left: Proportion of medium-difficulty validation levels solved vs.
 environment steps used in training.
-Curves show the RNN’s performance with a specific number of forced thinking steps at episode start, along with a ResNet baseline.
-Bottom: Estimated planning effect: the 8-steps minus the 0-steps curve.
-We see that reasoning emerges in the first 70M steps and keeps increasing for the hard levels, but decreases for medium levels.
+Curves show the RNN’s performance with a specific number of thinking steps (forced no-op actions) at episode start, along with a ResNet baseline.
+Bottom left: Estimated planning effect: the 8-steps minus the 0-steps curve.
+We see that reasoning emerges in the first 70M steps and keeps increasing for the hard levels (red), but decreases for medium levels (blue).
+Right: Linear probe predictions (arrows) for the direction the box will move in.
+Opacity is proportional to the number of steps in the episode in which the probe predicted a particular arrow.
+Correct predictions are in green and incorrect in red, the red predictions show other plausible plans for Box 4.
+This probe causally affects the actions of the agent, as described in Section 4.
+Boxes are numbered in the order in which they arrive at their final target.
 1 Introduction
-In many tasks, performance of both humans and sufficiently advanced neural networks (NNs) improves with more reasoning: whether giving a human more time to think before taking a chess move, or prompting a large language model to reason step by step (Kojima et al., 2022).
-Thus, understanding reasoning in such NNs is key to predicting their behavior in novel situations.
+In many tasks, the performance of both humans and sufficiently advanced neural networks (NNs) improves with more reasoning: whether by giving a human more time to think before taking a chess move, or by prompting or training a large language model (LLM) to reason step by step [Kojima et al., 2022, OpenAI, 2024].
 Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment.
-So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning (Hubinger et al., 2019) – may internalize goals different from that given by the training objective, leading to harmful misgeneralization (Di Langosco et al., 2022; Shah et al., 2022).
-This paper presents a 1.29M parameter recurrent neural network (RNN) that clearly benefits from extra time to reason.
-Following Guez et al.
-(2019), we train a convolutional LSTM to play Sokoban, a challenging puzzle game that remains a benchmark for planning algorithms (Peters et al., 2023) and reinforcement learning (Chung et al., 2024).
-We believe understanding this RNN is a useful first step towards understanding reasoning in sophisticated neural networks.
-The network is complex enough to be interesting: it reasons in a challenging environment and benefits from extra thinking time.
-At the same time, it is simple enough to be tractable to reverse engineer with current mechanistic interpretability techniques, owing to its small size and relatively straightforward behavior.
-We make three key contributions.
-First, we replicate and open-source an RNN whose performance improves with thinking time (Guez et al., 2019).
-Second, we analyze its training dynamics, finding that the thinking-time behavior emerges early in training and is incentivized by the setup (Figure 1, Sections 3 and 4.1).
-Third, we conduct a detailed empirical analysis of the network’s behavior.
-We quantitatively investigate how increased thinking changes network behavior (Section 3), showing that it reduces myopic actions (Figure 5).
-Furthermore, we present case studies of planning behavior that suggest the RNN naturally takes time to think (Section 4), which may explain the reduced planning effect on medium levels (Figure 1).
-2 Training the test subject
-We closely follow the setup from Guez et al.
-(2019), using the IMPALA (Espeholt et al., 2018) reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture, as well as a ResNet.
+So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning [Hubinger et al., 2019] – may internalize goals different from the training objective, leading to harmful misgeneralization [Di Langosco et al., 2022, Shah et al., 2022].
+Within goal-oriented reasoning, this work distinguishes between plans and search algorithms.
+A plan is a sequence of (perhaps abstract) actions which an agent intends to take.
+A search algorithm is an algorithm that considers many plans, evaluates them according to their predicted outcomes, and picks the plan evaluated as best.
+Search algorithms perform better the more plans they can evaluate and, with enough compute, generalize very well to novel problem instances [Russell and Norvig, 2009].
+This work investigates a recurrent neural network (RNN) that plays Sokoban [Guez et al., 2019], a challenging puzzle game that remains a benchmark for planning algorithms [Peters et al., 2023] and reinforcement learning [Chung et al., 2024].
+Using linear location-wise probes from concurrent work [Bush et al., 2024], we show that its activations represent a plan in a way which causes the RNN’s actions.
+The behavior of the RNN suggests that it also performs search.
+We find that the RNN sometimes takes actions which return to previous states of the environment.
+This overwhelmingly happens at the beginning of episodes and can be substituted by no-ops, suggesting the RNN deliberately goes in cycles to get more time to think.
+Giving the RNN time to think with no-ops [Guez et al., 2019] disproportionately helps with levels that are difficult (Figure 1), long, or that require waiting many steps before the first reward (Figure 3).
+Using insights from the plan probes and inspecting activations, we perform model surgery to uncover latent generalization capabilities of the RNN.
+The RNN is trained only on $10\times 10$ levels with 4 boxes (and its output MLP can only handle $10\times 10$ levels), but we are able to make it play on larger, nontrivial levels with more boxes (Figure 2).
+Altogether, evidence from probes and behavior show that the RNN causally represents a plan.
+The evidence highly suggests that the RNN implements an online search algorithm which scales with computation, although we cannot rule out multi-step heuristics.
+Consequently, we believe this RNN is an excellent model organism for understanding planning in sophisticated neural networks.
+Its behavior is complex enough to be interesting and its size (1.29M parameters) is small enough to be tractable to reverse engineer.
+This work shows that supervised probes and sparse autoencoders [Huben et al., 2023, Bricken et al., 2023, Gao et al., 2024] are not enough to fully do so, but it is likely that other mechanistic interpretability techniques such as circuit analysis [Olah et al., 2020, Geiger et al., 2021, Conmy et al., 2023] will reveal more about the algorithm which produces plans.
+Figure 2: The $\text{DRC}(3,3)$ is limited to $10\times 10$ grids by a fixed-dimension MLP.
+With model surgery, the ConvLSTM backbone generalizes to much more challenging settings than the $10\times 10$ training set grids.
+Left: XSokoban-31, which the $\text{DRC}(3,3)$ solves with 0 thinking steps.
+Right: XSokoban-29, which the $\text{DRC}(3,3)$ solves but only after 128 thinking steps.
+More examples in Appendix B.
+2 Setting up the test subject
+We train an agent closely following the setup from Guez et al.
+[2019], using the IMPALA [Espeholt et al., 2018] reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture.
+We also train a ResNet baseline.
 We open-source both the trained networks  [^1] https://github.com/AlignmentResearch/learned-planner and training code  [^2] https://github.com/AlignmentResearch/train-learned-planner .
+For more details about architectures and training, see Appendix A.
+$\text{DRC}(D,N)$ architecture.
+Most of this paper focuses on the behavior and representations of a $\text{DRC}(3,3)$ neural network [Guez et al., 2019].
+The core of this network is a $D$-layer ConvLSTM [Shi et al., 2015], which is repeatedly applied $N$ times per environment step.
+The output of the last ConvLSTM layer (the $D$th layer) is fed to the input at the next tick, effectively giving the network $D\cdot N$ layers of sequential computation to output the next action.
+We take $D=3,N=3$.
+A linear combination of the mean- and max-pooled activations of the ConvLSTM is also fed to the next step, thus letting it communicate quickly across the receptive field (known as pool-and-inject).
+An encoder block consisting of two $4\times 4$ convolutions processes the input, which is also fed to each ConvLSTM layer.
+At the end, an MLP with 256 hidden units transforms the flattened ConvLSTM outputs into the policy (actor) and value function (critic) heads.
 Dataset.
@@
 The player’s goal is to push all boxes onto target tiles while navigating walls.
-We use the Boxoban dataset (Guez et al., 2018), consisting of $10\times 10$ procedurally generated Sokoban levels, each with 4 boxes and targets.
+We use the Boxoban dataset [Guez et al., 2018], consisting of $10\times 10$ procedurally generated Sokoban levels, each with 4 boxes and targets.
 The edge tiles are always walls, so the playable area is $8\times 8$.
@@
 Guez et al.
-(2019) generated these by filtering levels that cannot be solved by progressively better-trained DRC networks, so easier sets occasionally contain difficult levels.
-In this paper, we use unfiltered-train (900k levels) to train networks.
-To evaluate them, we use unfiltered-test (1k levels) [^3] We prefer unfiltered-test rather than unfiltered-validation so the numbers are directly comparable to Guez et al.
-(2019)., medium-validation (50k levels), and hard ($\sim$3.4k levels), which do not overlap.
+[2019] generated these sets by filtering levels that cannot be solved by progressively better-trained DRC networks, so easier sets occasionally contain difficult levels.
+In this paper, we use the unfiltered-train (900k levels) set to train networks.
+To evaluate them, we use the unfiltered-test (1k levels) [^3] We use unfiltered-test rather than unfiltered-validation so the numbers are directly comparable to Guez et al.
+[2019]., medium-validation (50k levels), and hard ($\sim$3.4k levels) sets, which do not overlap.
+To test $\text{DRC}(3,3)$ generalization to different sizes, we use the levels collected by Þorsteinsson [2009] (see Appendix B).
 Environment.
 The observations are $10\times 10$ RGB images, normalized by dividing each component by 255.
-Each type of tile is represented by a pixel of a different color (Schrader, 2018).
-See Figure 6 for examples.
+Each type of tile is represented by a pixel of a different color [Schrader, 2018], an example is in Figure 12.
 The player can move in cardinal directions (Up, Down, Left, Right).
-The reward is -0.1 per step, 1 for putting a box on a target, -1 for removing a box, and 10 for finishing the level.
-To avoid strong time correlations during learning, each episode resets at a uniformly random length between 91 and 120 time steps.
-Figure 2: Left: A Sokoban level from the hard set. Right: the same level as the NN sees it, one pixel per tile. Walls are black, boxes are brown, targets are pink and the robot is green.
+The reward is -0.1 per step, 1 for placing a box on a target, -1 for taking it off, and 10 for finishing the level by placing all of the boxes.
+The time limit for evaluation is 120 steps, though large levels in Section 5.1 use 1000 steps.
+3 Behavioral evidence of search
+3.1 Effects of thinking time: non-myopic network
+First, we examine the effect of thinking time: introducing no-ops at the beginning of an episode, while letting the RNN process its hidden state.
+As in Guez et al.
+[2019], adding 6 extra thinking steps lets the $\text{DRC}(3,3)$ solve an extra $4.7\%$ of levels, slightly decreasing when going up to 16 steps (Figure 5, left).
+We find that the extra thinking time disproportionately helps with more difficult levels, as measured by longer optimal solutions (Figure 3, right).
+We also checked whether thinking time correlates with the number of nodes expanded by an A* search, but it does not (Figure 14).
+Thinking time helps varying amounts depending on training time and level difficulties: Figure 1 (bottom left) shows that most of the planning effect steeply appears during the first 70M steps of learning.
+Over the remaining training time, the RNN benefits more from thinking steps for the hard levels, but less for the medium levels; indicating perhaps that it learns better heuristics for them.
+We also find that thinking time disproportionately helps with levels in which the agent can greedily put a box to target early, but where doing so is detrimental to solving the level.
+Figure 3 (left, b) shows this: with 0 thinking steps, the first box is pushed to target on average 4 steps too early; compared to the solution reached at 6 thinking steps.
+Some of this is driven by the forced thinking steps preventing the $\text{DRC}(3,3)$ from taking a catastrophic action in the first few moves.
+However, the average time to first box is more than 6, even for 0 thinking steps, so there must also be many cases in which the 6 extra thinking steps give the NN time to find a better solution, rather than just preventing it from greedily pushing the first box.
+It must also be less greedy for boxes 2-3.
+Figure 3: Left: Average time step to place each box $B_{i}$ on target for different numbers of thinking steps.
+(a) Averages across all levels where the box $B_{i}$ is placed on target by DRC with $N$ thinking steps.
+(b) Averages for all levels solved by $6$ thinking steps but not solved by 00 thinking steps.
+More thinking steps makes the DRC avoid greedy strategies in favor of long-term return.
+Right: Average optimal solution length of levels grouped by the number of thinking steps at which the level is first solved.
+Levels that take longer to solve tend to be harder.
+NS stands for “not solved”.
+3.2 Agent “paces” to get more computation
+On occasion, the DRC exhibits a curious behavior: the agent “paces” in a cycle, returning to the same location, without touching any box.
+The per-step penalty makes this behavior naively sub-optimal, because it does not advance the puzzle state.
+Since the $\text{DRC}(3,3)$’s performance improves with thinking steps, could the agent be using pacing to gain computation time to improve its plan and so avoid irreversible actions which may block the solution?
+The following evidence suggests the agent has developed this meta-strategy:
+The agent paces mostly at the start of episodes.
+In some levels, a single sub-optimal step can lock the level, which makes calculating a valid plan in early steps beneficial.
+Figure 4 (left) shows most cycles start in the first 5 steps of the episode, and (middle) forcing the model to think for six steps eliminates about 75% of these early cycles.
+Thinking steps almost entirely substitute pacing
+We run the $\text{DRC}(3,3)$ on medium-validation levels and record when and where cycles occur. We merge all cycles which overlap with other cycles, for a total of 13 7021370213\,70213 702 cycles.
+We then make the DRC take $N$ thinking steps just before it would start a cycle of length $N$.
+As shown in Figure 4 (right), these thinking steps replace the cycles: comparing the cycle and the substitution, the DRC follows the same trajectory for at least $60\%$ of the levels and for a minimum of $30$ steps [^4] For context, the median solution length for train-unfiltered is exactly 30 steps..
+If our hypothesis that cycles are utilized for thinking holds true, we would expect to see uninterrupted periods of action after each cycle.
+In the $N$ steps after an $N$-step cycle concludes, $98.8\%$ of trajectories have no cycles.
+When substituting them for NOOPs, this becomes $82.4\%$, lower but still much higher than $0\%$; indicating that thinking steps largely remove the NN’s perceived need for cycles.
+Comparing $\text{DRC}(3,3)$ to ResNet
+It is possible that the DRC is pacing because it is not optimized enough.
+The ResNet is much worse than the $\text{DRC}(3,3)$ at Sokoban, but it is optimized enough to almost never perform cycles: since it has no memory, any cycles would make it stuck.
+So the amount of training is definitely enough to remove cycles if they have a big impact, and we believe it is likely to also remove cycles if they have a moderate negative impact.
+It could also be that the DRC is computing a multi-step heuristic (e.g.
+one that depends on the convolutional receptive field).
+Against this, the pool-and-inject layer in the DRC makes the receptive field less important for communication, and the ResNet (which performs worse) has a much larger receptive field per forward pass.
+Together, this is weak evidence that receptive-field heuristics are not the main mechanism.
+Heuristics may still play a role.
+For example, due to the -0.1 per-step penalty and discount rate, it is rational for the training process to make the RNN place boxes greedily (hence the increasing time-to-box in Figure 3, middle).
+It could be that the planning effect disappearing for medium levels (Figure 1, left) is due to better heuristics about when to be greedy and when to take time to think.
+Figure 4: Left: Histogram of cycle start times on the medium-difficulty validation levels.
+Middle: Total number of cycles the agent takes in the first 5 steps across all episodes in medium-difficulty validation levels with $N$ initial thinking steps.
+Right: on all levels, we replace $N$-length cycles with $N$ thinking steps and check the proportion of trajectories which are equal between these two treatments, for $x$ steps after the cycle.
+4 The $\text{DRC}(3,3)$ causally represents its plan
+In order to improve the plan with search, as the previous section suggests, the agent must have some internal representation of the plan.
+The absence of such representation would force us to discard the theory that the agent is doing search.
+This section presents probes which predict the future actions of the $\text{DRC}(3,3)$ and other various future features of the environment.
+Of these, the one which predicts the future direction boxes will move in has a strong causal effect in the actions of the RNN.
+Following the standard of evidence in Li et al.
+[2023], we declare this conclusive evidence that the $\text{DRC}(3,3)$ represents and uses plans.
+The spatial structure of the probes and some of the targets are due to Bush et al.
+[2024].
+We also observe the plan grow during thinking steps and, by intervening on the probes, are able to lock the RNN into a plan when there are two options of equal value.
+However in most cases, if we stop intervening on the RNN activations with a suboptimal plan, the RNN comes up with a better plan online and follows it.
+We have some evidence for the $\text{DRC}(3,3)$ considering multiple plans: for example, in the level in Figure 1 (right) the box probe first considers moving box 4 down and right (shown in red), but then rethinks that and ends up taking different actions.
+However, we have been unable to find multiple simultaneous representations of plans, or anything that would decide between them such as the value of the plans.
+In all, this section provides very little evidence in favor of search, but conclusive evidence for a causal plan.
+Figure 5: Success rate of the RNN, and plan quality as measured by box-direction probes (Section 4.1), all increase with thinking steps.
+We measure plan quality by 1) summing the length of chains of probe-predicted directions, which start in boxes, and 2) simply counting the number of squares in which the box-direction probe predicts something.
+The F1 score of the plan probe also increases over time, suggesting that the DRC builds the plan in that time.
+The right plot also includes plan quality evaluation on the three sequential ticks per environment step of the DRC(3, 3) network.
+4.1 Probe methodology: training, intervention and targets
+Training.
+We train logistic or linear regression probes with L1 decay to predict various features of the environment from the agent’s activations.
+We train on two types of inputs.
+1) Grid-wise inputs [Bush et al., 2024]: each square in the $10\times 10$ grid is a different data point, potentially with a different label depending on location and timestep.
+The input is the $64$-dimensional LSTM state $(h,c)$ at a square, usually concatenated for the 3 layers.
+2) Global inputs: Each time step produces only one data point.
+The probes take as input the 64 channels of all of the squares in the 10x10 grid, resulting in a 6400-dimensional input for each layer.
+In both cases, the per-layer activations are concatenated for probes that are trained on multiple layers.
+The dataset consists of states collected by evaluating the $\text{DRC}(3,3)$ on the hard Boxoban levels, excluding the first 5 steps of each episode because the plan is still forming.
+For multi-class probes, training and F1 scores are computed as one-vs-all: the presence of absence of a particular class is the probe label.
+We search the best learning rate and L1 decay with grid-search by evaluating the F1 on a validation split of $20\%$ of timesteps from the hard levels.
+We take the first 1000 medium-validation levels as our test set to report the results (Table 1).
+Causal intervention.
+Given a probe that is predictive of future behavior, we edit the agent’s activations such that the probe predicts the behavior we desire to induce.
+To the extent that the agent carries out the edited behavior, we can conclude that it executes a plan encoded in the linear direction of the probe.
+For a linear probe on the activations $h$ with parameter vector $p$, we intervene in order to increase the logits for a particular behavior, given by $p\cdot h$ (we omit a scalar intercept for brevity).
+Unlike previous work in steering vectors for LLMs [Turner et al., 2023, Rimsky et al., 2023, Li et al., 2024], we found that simply adding a perturbation equal to a fixed multiple of $p$ caused the policy to degenerate and choose random actions.
+We instead add an adaptively-scaled multiple of $p$ to the activations, chosen to minimize the change in the activations while still causing the probe’s logit to be greater than a desired amount $\alpha$.
+More precisely, we find $\operatorname*{arg\,min}_{h^{\prime}}\|h^{\prime}-h\|_{2}^{2}$, with the constraint that $h^{\prime}\cdot p>\alpha$.
+The solution to this is $h^{\prime}=h+{\hat{p}}\max(0,\alpha-(h\cdot p))$, where ${\hat{p}}=p/\|p\|^{2}_{2}$.
+Targets.
+We choose targets that plausibly encode the steps of a plan.
+Most of these labels are for grid-wise probes, but the pacing and value probes are global.
+The direction probes are very similar to Bush et al.
+[2024].
+  • Agent-Directions probe: This probe has 5 outputs: NV (No Visit), UP, DOWN, LEFT, RIGHT.
+The probe takes as input the 64-dim hidden state activations at a square $(x,y)$ at timestep $t_{i}$.
+If the agent visits the square $(x,y)$ at a future timestep $t>t_{i}$, then the probe predicts the corresponding direction that the agent takes from that square.
+For squares that are visited multiple times in the future, the direction corresponding to the first future visit is taken as the target.
+If a square is never visited in a future timestep, the probe predicts NV.
+There is one single multinomial probe, for all time steps and locations $(x,y)$.
+  • Boxes-Directions probe: This probe is similar to the above probe, except that it is trained to predict the direction in which any of the four boxes will move at a given square in the future.
+The rationale for this probe is that the boxes, and constraints on their movement given by obstacles, are the main difficulty with Sokoban puzzles.
+  • Next-Box probe: Predicts $1$ on the square of the box that the agent will move next and 00 for every other square.
+  • Next-Target probe: Predicts $1$ on the square of the target that the agent will put a box in next, and 00 for every other square.
+  • Pacing probe: The global label is 1 if the agent is currently in a cycle, and 0 otherwise.
+  • Value probe: The global label is the numerical value that the critic head outputs.
+[…]
+Probe target Predictive Probe Results (a) Causal Probe Results (b)
+Sparse Probes Causal Effect
+Best F1 F1 $L_{0}$-norm $\alpha$ Average Best-case
+Box-directions $86.4\pm 0.1$ $73.1\pm 0.1$ $63$ 30 $43.7\pm 0.6$ $77.8\pm 0.9$
+Agent-directions $72.3\pm 0.1$ $61.3\pm 0.1$ $121$ 10 $7.1\pm 0.3$ $20.7\pm 0.7$
+Next box $74.2\pm 0.4$ $69.7\pm 0.5$ $51$ 40 $5.5\pm 1.0$ $15.1\pm 2.5$
+Next target $54.3\pm 0.5$ $44.0\pm 0.5$ $32$ 30 $4.6\pm 0.8$ $13.2\pm 2.0$
+Pacing $31.0\pm 1.8$ $31.0\pm 1.5$ $5$
+4.2 Probe evaluation and causality
+Probe predictive power.
+Most probes are quite predictive, as seen in Table 1.
+The exception is the pacing probe, with $F_{1}=31.0\%$, which is not much better than the constant 1 probe, which has $F_{1}=12.8\%$.
+This leads us to conclude that the $\text{DRC}(3,3)$ does not represent whether it is in a cycle or not.
+See Figure 1 (right) and Figure 6 for visualizations of agent-direction and box-direction probes [^5] The supplementary material provides visualization videos of all the probes across several levels..
+For the value function probe, we compute the fraction of variance explained $R^{2}$.
+If we train a global probe, $R^{2}$ is very high: $97.7\%$$99.7\%$ depending on the layer.
+However, grid-wise probes obtain much worse but still passable results: $41.0\%$$79.2\%$ depending on layer.
+We visually checked whether the grid-wise probe reads off the values of different plans, but could not find any such pattern.
+Almost all the performance of the global probe is recovered by training on the mean-pooled inputs: $95.2\%$$99.5\%$.
+It is likely that the global and mean-pooled value probes are indirectly counting the number of squares the agent will step on, which almost fully determines the value, and we know is possible due to the future-direction probes.
+Only the Box-Directions probe is strongly causal.
+We measure how causal the probes are by intervening at every step on the hidden state $h,c$ at the layer the probe was trained on, following the procedure in section 4.1 and checking whether the agent follows that action.
+We report the result in Table 1(b) by performing a grid-search over the strength of intervention $\alpha$ and picking the best value.
+Despite our adaptive scaling of the intervention, a high value of $\alpha$ can still disturb the agent such that it starts performing random actions whereas a low value of $\alpha$ may not be sufficient to cause the agent to follow the intended behavior.
+Most of the probes are not causal: only the “boxes directions” probe (used in Figure 1) is causal while the “agent-direction” probe mildly causal.
+Even then, we have observed that it can only change boxes directions when they would not deadlock the level in a naive way.
+That is, if you try to make the model push a box onto a wall, such that the box would not be able to reach a target anymore, it does not.
+For this reason, we introduce the Best-case causal effect in Table 1(b): try all three directions that are not the actual direction that the box follows, and count the probe as “causal” if it works for any of them (e.g.
+because they do not naively deadlock the level).
+The box-direction probe’s coverage increases during thinking time.
+Figure 5 (middle) show that the amount of squares in which the probe predicts a direction, be they lone or linked to a box, grows over thinking steps.
+Its description of what the box movements will be also greatly improves after a few steps of thinking Figure 5 (right).
+4.3 Bistable and unstable plans in toy environments
+The causal intervention results from Table 1 show that the box-directions probe is much more causal than the agent-directions probe.
+We further study this in an isolated setting.
+Figure 6 (left) shows a level with only a single box next to a target, where the agent has only two equally good paths it can follow.
+Without intervening, the agent first takes the path on the right by moving three steps, while the agent-direction probe grows towards the box.
+After the first three moves, the chain connects.
+The agent then takes a U-turn to follow the path on left.
+This is very unlike the optimal behavior, which is to commit to any path and follow it.
+We can get the agent to commit to one path early by intervening with the agent-direction probe on any of the two paths on the first step only, as shown in Figure 6 (middle).
+When we construct a similar level where there are two equally good paths for a box to follow, we find that the agent-direction probe is no longer causal; but the boxes direction probe does get the agent to commit.
+Figure 6 (right) shows an almost empty level.
+When we intervene with the box-directions probe on the first four steps, the agent follows the laid out path in those steps, but then quickly switches to a more optimal path with fewer turns.
+The laid out path (not taken, shown in red) is not followed after stopping the intervention.
+This analysis provides strong evidence that the network primarily builds upon the boxes directions represented in the activations and relies on the agent-direction when the boxes directions do not inform the subsequent actions.
+Figure 6: Left: A custom level where the agent has two equally good paths to follow.
+The arrows show the prediction of the agent-direction probe with opacity proportional to the number of times an arrow was predicted across all the steps.
+The green and red arrows are correct and incorrect predictions, respectively.
+The agent behaves suboptimally by returning to the start and going left after first going right for three steps.
+Middle: The agent takes the path on the right after intervening with the corresponding arrows using the agent-direction probe on the first step.
+The same happens on the left if we intervene on that path without the suboptimal steps right of the undisturbed agent.
+Right: An empty level with box-directions probe intervening on the first four steps which are followed correctly by the agent on those four steps.
+When the intervention is removed on the fifth step, the agent computes the simpler path in green and doesn’t follow the path laid out earlier (in red).
+5 Model surgery exposes latent generalization capabilities
+A common feature of search algorithms is that they generalize very well to unusual problems in their domain.
+This was the main thrust of the argument of Guez et al.
+[2019]: they exhibited many ways in which the $\text{DRC}(3,3)$ generalized to levels outside of the training distribution to argue that it does search.
+5.1 Generalizing beyond $10\times 10$ inputs and training examples
+This section goes beyond the capabilities of $\text{DRC}(3,3)$ as a black-box unit.
+By looking for interpretable features in the activations of the ConvLSTM core, we found that layer 3, at the last tick, has four channels which represent the next action to take.
+Since the network up until the ConvLSTM core is completely convolutional, we can evaluate it on inputs of any size – unlike $\text{DRC}(3,3)$ as a whole, which only works on $10\times 10$ inputs, and is trained only on levels with four boxes.
+Spatial aggregation.
+We start with the Layer 3 grid-wise next-action probes from Table 3.
+For each spatial location, these predict the action the $\text{DRC}(3,3)$ will take in this step.
+We attempt three different spatial aggregation methods: mean-pooling, max-pooling, and the proportion of locations with positive probe readouts for a particular action.
+We learn their relative weights and a baseline for each action by using Adam with learning rate $\ell=10^{-3}$, annealed linearly to 0 for 10000 steps.
+Table 2 lists the learned parameters, which place most weight on the mean-pooling aggregation rule.
+The training set consists of 3000 levels, 1000 from each of the Boxoban unfiltered, medium, and hard training sets.
+Table 2: Weight and bias for transforming direction probes to predictions
+Mean Max Positive proportion
+Weight 1.2086 -0.0582 0.2070
+Up Down Left Right
+Bias 0.3337 -0.0921 -0.0632 -0.0539
+Results.
+The aggregated features obtain 83% accuracy at predicting the MLP’s action output on the training set, which is less than expected given the $90\%+$ F1 scores of probes in Table 3.
+With only seven parameters, we did not bother collecting an i.i.d.
+test set and went straight to the generalization tests.
+The adjusted ConvLSTM is able to solve many levels that are out-of-distribution: larger than $10\times 10$ in both dimensions, and with more than four boxes, e.g.
+those in Figure 2.
+Out of our test set collected from Þorsteinsson [2009], it can solve $59/480$ levels with both dimensions $>10$, $30/203$ levels with only one dimension $>10$ and $180/482$ levels with both dimensions $\leq 10$.
+Appendix B has a breakdown by level collection.
+5.2 Looking for interpretable features with Sparse Autoencoders
+The way we found out actions were represented in the ConvLSTM was with sparse autoencoders (SAEs).
+We inspected an SAE [Huben et al., 2023, Bricken et al., 2023] trained on the last layer with the top-k activation [Gao et al., 2024, Tamkin et al., 2024], $k=8$.
+However, we then noticed that the action features were also represented directly in the ConvLSTM channels.
+Table 3 compares raw channels, SAE features, and linear probes for how well they match the interpretation of “the feature fires when the $\text{DRC}(3,3)$ takes a particular action”.
+We find that the raw channels and linear probes are similarly good, and both are considerably better than the SAE.
+Thus, for these features, SAEs actually worsened interpretability. We also found three other interpretable features in the SAE, which were also represented directly in the channels (Appendix E).
+Table 3: Action features scores across channels, probes, and SAE features
+Feature Channel SAE Feature Probe
+Number Prec Rec F1 Number Prec Rec F1 Prec Rec F1
+Up L3C29 95.7 88.1 91.7 L3F270 93.9 76.2 84.1 97.5 86.5 91.7
+Down L3C8 98.4 80.8 88.8 L3F187 98.0 79.1 87.6 97.6 86.9 91.9
+Left L3C27 85.5 84.6 85.1 L3F244 96.1 63.2 76.2 83.5 86.6 85.0
+Right L3C3 97.0 86.9 91.7 L3F385 94.6 78.5 85.8 97.6 87.4 92.2
+6 Related work
+Interpreting agents and planning.
+Several works have attempted to find the mechanism by which a simple neural network does planning in mazes [Mini et al., 2023, Knutson et al., 2024, Brinkmann et al., 2024], gridworlds [Bloom and Colognese, 2023], and graph search [Ivanitskiy et al., 2023].
+Men et al.
+[2024] investigates LLM reasoning in a simple block-stacking task.
+We believe the DRC we present is a clearer example of an agent than what these works focus on, and should be similarly possible to interpret.
+Concurrent work [Bush et al., 2024] interprets another DRC Sokoban agent and can also predict future actions. That team invented grid-wise probes and the Agent-Directions and Box-Directions targets.
+Other works interpret superhuman game-playing agents.
+Jenner et al.
+[2024] find evidence of lookahead in Leela Chess Zero [Team, 2018], without predicting future actions.
+McGrath et al.
+[2021], Schut et al.
+[2023] find interpretable concepts in AlphaZero [Silver et al., 2017].
+Probing for world models.
+Planning usually requires a model of the world.
+Li et al.
+[2023], Nanda et al.
+[2023b], Karvonen [2024] examine agents trained to predict possible moves in games, and are able to probe for the board state.
+Wijmans et al.
+[2023] also probe the state of navigation neural networks.
+Gurnee and Tegmark [2023] find a representation of positions in Earth for location-associated tokens in an LLM.
+Goal misgeneralization and mesa-optimizers for alignment.
+From the alignment perspective, AIs optimizing monomaniacally for a goal have long been a concern (e.g.
+Yudkowsky, 2006, Omohundro, 2008; see the preface of Russell, 2019).
+In a machine learning paradigm [Hubinger et al., 2019], the goal of the training system is not necessarily optimized; instead, the NN may optimize for a related or different goal [Di Langosco et al., 2022, Shah et al., 2022] or for no goal at all.
+Chain-of-thought faithfulness.
+Large language models use chain of thought, but are they faithful to it, or do they think about their future actions in other ways [Lanham et al., 2023, Pfau et al., 2024]?
+One could hope that LLMs perform all long-term reasoning in plain English, allowing unintended human consequences to be easily monitored, as in Scheurer et al.
+[2023].
+Ethical treatment of AIs.
+Do AIs deserve moral consideration?
+Schwitzgebel and Garza [2015] argue that very human-like AIs are conceivable and clearly deserve rights.
+Tomasik [2015] suggests that most AIs deserve at least a little consideration, like biological organisms of any species [Singer, 2004].
+But what does it mean to treat an AI ethically?
+Daswani and Leike [2015] argue that the way to measure pleasure and pain in a reinforcement learner is not by its absolute amount of return, but rather by the temporal difference (TD) error: the difference between its expectations and the actual return it obtained.
+If the internals of the NN have a potentially different objective [Hubinger et al., 2019, Di Langosco et al., 2022], then the TD error should come from a place other than the critic head.
+Research into learned search algorithms is an early step toward finding the learned-reward internal TD error, if it exists.
+This could be a higher-assurance complement to simply asking the AI [Perez and Long, 2023].
+Fully reverse engineering small networks.
+Many works reverse engineer all of, or most of, a small NN that does an algorithmic task [Nanda et al., 2023a, Chughtai et al., 2023, Zhong et al., 2023, Quirke and Barez, 2023].
+Neural network architectures that reason.
+Many papers try to enhance NN thinking by altering the training setup or architecture.
+Schwarzschild et al.
+[2021b, a], Bansal et al.
+[2022] improve and evaluate a specific kind of RNN training that produces algorithmic thinking, though Knutson et al.
+[2024] argue that they do not generalize enough and thus do not implement the correct algorithm.
+Other works endow RNNs with variable amounts of computation per step, with [Chung et al., 2024] or without [Graves, 2016] an explicit world model.
+7 Conclusion
+We replicated a small RL-based Sokoban agent, whose performance benefits from increased compute in the form of externally imposed no-op steps [Guez et al., 2019].
+We find that the agent learns to explicitly exploit this capability by pacing, which substitutes for the external no-op thinking steps.
+Extra computation helps with longer levels and those that require non-myopic thinking.
+We show that the agent causally represents plans (sequences of actions to take).
+The most fundamental plan representation we find is the box-direction probes, which can predict and control where the agent pushes a box.
+By building up a chain of box probe steps over time, the agent is able to figure out how to place boxes to solve the level.
+Secondarily, we find agent-direction probes that take lower precedence than the box-direction probes, but can control the agent moves in isolated setups.
+We additionally find that the box probe chains get longer and longer with more thinking steps (Figure 5).
+Finally, we find that extracting a decision from the plan takes place almost entirely within the ConvLSTM, with the MLP doing little more than decoding it.
+To demonstrate this, we show that the agent can generalize to levels bigger than what it had ever seen during training by manually replacing its MLP decoder based on our mechanistic understanding.
+We hope the model organism we trained and open-source will catalyze research into NNs that learn to plan, and to understand their inner objectives.
+Acknowledgements
+We would like to thank ChengCheng Tan and Siao Si Loo for editing this paper, and the rest of the FAR team for general support.
+We also thank Alex F.
+Spies and anonymous reviewers for improving the paper with their comments and edits.
+We are also very grateful for the many questions about their RL setup that Arthur Guez and Timothy Lillicrap answered, as well as inspiration from Stephen Chung and Scott Goodfriend.
+Reproducibility Statement
+We provide the code and the trained models in the supplementary materials.
+We open-source our code and trained model(s).
+Additionally, sections 2, 4.1 and appendices A, D, and E describe the specific experimental details and settings we used to carry out our experiments.
+Appendix A Training the test subject
 $\text{DRC}(D,N)$ architecture.
 Guez et al.
-(2019) introduced the Deep Repeating ConvLSTM (DRC), whose core consists of $D$ convolutional LSTM layers with 32 channels and $3\times 3$ filters, each applied $N$ times per time step.
+[2019] introduced the Deep Repeating ConvLSTM (DRC), whose core consists of $D$ convolutional LSTM layers with 32 channels and $3\times 3$ filters, each applied $N$ times per time step.
 Our $\text{DRC}(3,3)$ – or just DRC for brevity – has $1.29$M parameters.
 Before the LSTM core, two convolutional layers (without nonlinearity) encode the observation with $4\times 4$ filters.
-The LSTM core uses $3\times 3$ convolutional filters, and a nonstandard tanh on the output gate (Jozefowicz et al., 2015).
-Unlike the original ConvLSTM (Shi et al., 2015), the input to each layer of a DRC consists of several concatenated components:
+The LSTM core uses $3\times 3$ convolutional filters, and a nonstandard tanh on the output gate [Jozefowicz et al., 2015].
+Unlike the original ConvLSTM [Shi et al., 2015], the input to each layer of a DRC consists of several concatenated components:
   • The encoded observation is fed into each layer.
   • To allow spatial information to travel fast in the ConvLSTM layers, we apply pool-and-inject by max- and mean-pooling the previous step’s hidden state.
-We linearly combine these values channel-wise before feeding them as an input to the next step.
+We linearly combine these values channel-wise before feeding them as input to the next step.
   • To avoid convolution edge effects from disrupting the LSTM dynamics, we feed in a $12\times 12$ channel with zeros on the inside and ones on the boundary.
@@
 This is a convolutional residual neural network, also from Guez et al.
-(2019).
+[2019].
 It serves as a non-recurrent baseline that can only think during the forward pass (no ability to think for extra steps) but is nevertheless good at the game.
@@
 The first two blocks have 32 channels, and the others have 64.
-Each block consists of a convolution, then two (relu, conv) sub-blocks, which each split off and are added back to the trunk.
+Each block consists of a convolution, followed by two (relu, conv) sub-blocks, each of which splits off and is added back to the trunk.
 The ResNet has $3.07$M parameters.
@@
 [^7] The A* solutions may be of independent interest, so we make them available at https://huggingface.co/datasets/AlignmentResearch/boxoban-astar-solutions/.
-3 Quantitative behavior analysis
-Figure 3: Proportion of medium-difficulty validation levels solved by DRC with $x$ time steps to think.
-The dotted line shows the ResNet’s solve proportion.
-The proportion of solved levels increases with thinking time, but only up to 6 steps.
-3.1 Recurrent networks are better at Sokoban
-After training both the recurrent $\text{DRC}(3,3)$ network and feed-forward ResNet for 2B environment transitions, the $\text{DRC}(3,3)$ network is able to solve $99.3\%$ of unfiltered test levels, while the ResNet solves only $97.9\%$.
-This gap is particularly notable given that the ResNet has more than double the parameters of the $\text{DRC}(3,3)$ network, so the ResNet has more model capacity and sequential thinking time.
-These results are consistent with Guez et al.
-(2019), see Appendix B.
-The performance gap widens when considering just the hard levels, of which the $\text{DRC}(3,3)$ network solves $42.8\%$ and the ResNet $26.2\%$.
-Counterintuitively, we find that – controlling for performance on the unfiltered test set – recurrent networks are not comparatively better at harder levels than the ResNet.
-Figure 10 plots the performance of model checkpoints on the unfiltered test set against medium-difficulty and hard levels.
-We find that both model architectures follow the same curve: the recurrent $\text{DRC}(3,3)$ network is further to the right, but follows the same trend.
-3.2 Thinking time improves performance
-We find the performance of our trained $\text{DRC}(3,3)$ improves given additional thinking time, replicating Guez et al.
-(2019).
-We evaluate the DRC network on all medium-difficulty validation levels.
-To induce thinking time, we repeat the initial environment observation $n$ times while advancing the DRC hidden state, then let the DRC policy act normally.
-Figure 3 shows the DRC’s performance for $n\in\{0,1,2,4,6,8,10,12,16\}$ steps, alongside the ResNet baseline.
-The DRC’s success rate improves from $76.6\%$ to $81.3\%$ with more thinking steps, with peak performance at 6 steps.
-The effect ($4.7\%$ difference) is comparable to that observed by Guez et al.
-($4.5\%$).
-Figure 4: Average optimal solution length of levels grouped by the number of thinking steps at which the level is first solved.
-Each of the 50,000 medium-difficulty validation levels is assigned to a group.
-Optimal solution length is a proxy for difficulty, and we can see that levels which take longer to solve tend to be harder.
-3.3 Planning emerges early in training
-Figure 1 shows the performance of the DRC on the medium-difficulty validation levels by training step.
-We see the DRC rapidly acquires a strong planning effect by 70M environment steps.
-This effect strengthens on the hard levels as training progresses, but slowly weakens for medium-difficulty levels.
-This suggests that the training setup and loss function reinforce the planning effect for harder levels and disincentivize it for levels that are comparatively easier.
-3.4 Do difficult levels benefit more from thinking?
-The A* solver gives us two proxies for the difficulty of a level: the length of the optimal solution, and the number of nodes it had to expand to find a solution.
-Figure 4 shows a positive correlation between the A* optimal solution length ($y$-axis) and the number of DRC thinking steps $n$ needed to solve a level ($x$-axis).
-In other words, taking extra more thinking time helps solve harder levels.
-On the other hand, Figure 11 (Appendix D) finds only a weakly positive correlation between the expanded nodes proxy and required thinking steps $n$, with results noisy for smaller $n$.
-We speculate that the ‘expanded nodes’ proxy metric is only weakly related to required $n$ due to the DRC having different heuristics than the one we use for A*, whereas the optimal solution length does not depend on heuristics.
-Table 1: Levels at 6 thinking steps.
-Level categorization Percentage
-Solved, previously unsolved 6.87
-Unsolved, previously solved 2.23
-Solved, with better returns 18.98
-Solved, with the same returns 50.16
-Solved, with worse returns 5.26
-Unsolved, with same or better returns 15.14
-Unsolved, with worse returns 1.36
-3.5 Other consequences of DRC thinking time
-While Figure 3 shows that the DRC can solve many more levels when given thinking steps at the beginning of an episode, in this section, we explore how the network’s behavior changes with thinking time, and if this has significant effects besides solving more levels.
-Thinking makes the DRC “patient.”
-One hypothesis is that, without the extra thinking steps, the DRC sometimes comes up with a greedy solution for some of the boxes and executes it right away.
-This puts some boxes into the wrong targets, and makes the puzzle unsolvable.
-Since the training signal penalizes every step it takes to solve the puzzle, it could be optimal for a bounded rational agent to avoid thinking and be “impatient” in this way: the agent may accumulate more return on average by being “impatient” and quickly solving most levels even if it sometimes errs by taking actions that make it impossible to solve other levels.
-For $n$ thinking steps, Figure 5 plots the average number of steps (time-to-box) until each of 4 boxes gets on a target, in the order in which they do so.
-The results support the hypothesis: levels that are solved at 6 thinking steps but unsolved at 0 thinking steps show increased time-to-box for boxes 1-3 but decreased time for box 4 (Figure 5(b)).
-In contrast, averaged across all levels times go down for all boxes with increased thinking (Figure 5(a)).
-Thinking often improves solutions for already-solved levels, but is sometimes harmful.
-Table 1 shows that an additional 6 thinking steps solves 6.87% of previously unsolved levels – but at the cost of no longer being able to solve 2.23% of previously solved levels, for a total improved solution rate of 4.64%.
-In addition to improving the solution rate, additional thinking improves returns on average by shortening solution length in 18.98% of episodes, consistent with the decreased timesteps to place boxes shown in Figure 5(a).
-However, the effect is inconsistent: most episodes show no change in returns, with returns even decreasing in 5.26% of episodes.
-Overall, extra thinking steps are not always beneficial but, on balance, they let the DRC solve more levels or improve return on already solved levels.
-Figure 5: Average time step to place each box $Bi$ on target for different thinking steps.
-(a) Averages across all levels where the box $Bi$ is placed on target by DRC with $n$ thinking steps.
-If box $Bi$ is not placed (on unsolved levels) we ignore the box, but we keep the other boxes in the level.
-(b) Averages for all levels solved by $6$ thinking steps but not solved by 00 thinking steps.
-(a) file 0, level 18
-[…]
-(b) file 0, level 53
-[…]
-(c) file 0, level 153
-[…]
-Figure 6: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps.
-Colors are as in Figure 2.
-Boxes and targets are paired in upper- and lower-case letters respectively, such that the DRC’s best solution places boxes on targets in alphabetical order.
-Videos available at this https URL; see Section 4 for behavior descriptions.
-Levels solved faster obtain higher return as the agent incurs the per-step penalty fewer times.
-4 Case studies: how does thinking change behavior?
-To gain a deeper understanding of the DRC networks’ behavior, we examine an example from each of three categories from Table 1: a solved, previously unsolved level; a level on which the DRC improved its return with extra thinking (solved faster); and a level on which the return worsens (solved slower).
-We select an illustrative example of each category from the first ten examples by level index.
-Thinking lets the DRC solve: Figure 6(a).
-In the no-thinking condition, the DRC first pushes box $C$ one square to the right.
-It then goes back to push $A$ to $a$, but it is too late: it is now impossible to push box $B$ onto $b$.
-The network pushes $C$ onto $c$, then $D$ onto $d$ and then remains mostly still.
-In contrast, after thinking, the DRC pushes $A$ to $a$ first, which lets it solve the level.
-Thinking speeds up solving: Figure 6(b).
-In the no-thinking condition, the DRC spends many steps going back and forth before pushing any boxes.
-First it goes down to $y$, then up to $c$, then down onto $z$, back up to $y$ and to $z$ again.
-It then proceeds to solve the rest of the puzzle correctly: push box $A$ onto $a$, prepare box $B$ on $x$ and box $C$ where $B$ originally was, push in boxes $B,C$ and finally $D$.
-In the thinking condition, the DRC makes a beeline for $A$ and then plays the exact same solution.
-Thinking slows down solving: Figure 6(c).
-In the no-thinking condition, the DRC starts by pushing box $C$ into position $y$, then pushes boxes $A,B$.
-On the way back down, ther DRC pushes $C$ onto $c$ and finally $D$ onto $d$.
-In contrast, after thinking for a bit, the DRC goes the other way and starts by pushing $B$ onto $b$.
-The solution is the same after that: $A$, $C$, then $D$; but the NN has wasted time trekking back from $B$ to $A$.
-4.1 Hypothesis: the DRC ‘paces’ to get thinking time
-The second case study above found the DRC sometimes moves around in cycles without touching any box.
-One hypothesis for why this occurs is that the DRC is using the extra steps to come up with a plan to solve the episode.
-To test this, we can check whether cycles in the game state occur more frequently near the start of the episode, and whether deliberately giving the DRC thinking time makes it stop going in cycles.
-Consistent with this hypothesis, Figure 7 (left) shows the majority of cycles start within the first 5 steps of the episode, and (right) forcing it to think for six steps makes about 75% of these cycles disappear.
-The episode start is not the only time at which the DRC moves in cycles.
-Figure 9 shows that replacing an $n$-length cycle with $n$ thinking steps leads the NN to have the exact same behavior for at least 60% of levels, for at least 30 steps after the cycle.
-For context, the median solution length for train-unfiltered is exactly 30 steps.
-Additionally, in $82.39\%$ of cases, doing this prevents cycles from starting in the $n$ steps after the thinking-steps conclude (Appendix C).
-Why does this behavior emerge during training?
-Thinking is useful for achieving higher return, so it should be reinforced.
-But it also has a cost, -0.1 per step, so it should be discouraged in easy levels that do not need the computation.
-We speculate that, as training advances and heuristics get tuned, the DRC needs to think in fewer levels, and it is better at knowing when it needs to pace.
-This would also explain why, after many training steps, the DRC network begins to benefit less from additional thinking steps at the start of the episode in Figure 1: the DRC finds levels easier and also knows to pace when thinking is needed.
-Figure 7: Left: Histogram of cycle start times on the medium-difficulty validation levels.
-Right: Total number of cycles the agent takes in the first 5 steps across all episodes in medium-difficulty validation levels with $n$ initial thinking steps.
-5 Related work
-Interpretability of agents and world models.
-Several works have attempted to find the mechanism by which a simple neural network does planning in mazes (Ivanitskiy et al., 2023; Mini et al., 2023), gridworlds (Bloom & Colognese, 2023), and graph search (Ivanitskiy et al., 2023).
-We believe the DRC we present is a clearer example of an agent than what these works focus on, and should be similarly possible to interpret.
-Other works have found emergent world models in sequence-prediction (Li et al., 2023) and navigation (Wijmans et al., 2023) neural networks.
-Goal misgeneralization and mesa-optimizers for alignment.
-From the alignment perspective, AIs optimizing monomaniacally for a goal have long been a concern (e.g.
-Yudkowsky, 2006; Omohundro, 2008; see the preface of Russell, 2019).
-In a machine learning paradigm (Hubinger et al., 2019), the goal of the training system is not necessarily optimized; instead, the NN may optimize for a related or different goal (Di Langosco et al., 2022; Shah et al., 2022), or for no goal at all.
-Chain-of-thought faithfulness.
-Large language models (LLMs) use chain of thought, but are they faithful to it or do they think about their future actions in other ways (Lanham et al., 2023; Pfau et al., 2024)?
-One could hope that LLMs perform all long-term reasoning in plain English, allowing unintended human consequences to be easily monitored, as in Scheurer et al.
-(2023).
-Reasoning neural network architectures.
-Many papers try to enhance NN thinking by altering the training setup or architecture (Bansal et al., 2022; Graves, 2016; Chung et al., 2024).
-Ethical treatment of AIs.
-Do AIs deserve moral consideration?
-Schwitzgebel & Garza (2015) argue that very human-like AIs are conceivable and clearly deserve rights.
-Tomasik (2015) suggests that most AIs deserve at least a little consideration, like biological organisms of any species (Singer, 2004).
-But what does it mean to treat an AI ethically?
-Daswani & Leike (2015) argue that the way to measure pleasure and pain in a reinforcement learner is not by its absolute amount of return, but rather by the temporal difference (TD) error: the difference between its expectations and the actual return it obtained.
-If the internals of the NN have a potentially different objective (Hubinger et al., 2019; Di Langosco et al., 2022), then the TD error should come from a place other than the critic.
-This paper is an early step toward finding the learned-reward internal TD error, if it exists.
-6 Conclusion
-We have developed and open-sourced a model organism for understanding reasoning in neural networks.
-In addition, we contribute a detailed empirical analysis of this network’s behavior.
-We believe this work will prove useful to the interpretability, alignment, and ethical treatment of AIs.
-First, we have trained a DRC recurrent network that plays Sokoban and benefits from additional thinking steps at test time, replicating Guez et al. (2019).
-Second, we have shown that thinking for longer helps solve harder levels, and makes the DRC better at levels that require longer-sighted behavior.
-Without intervention, thinking sometimes takes the form of the DRC ‘pacing’ at the beginning or middle of the episode, in a way which can be substituted by repeating the same input.
-This suggests the network is deliberately using more computation by pacing.
-Finally, we have shown that the training setup incentivizes the planning effect at the start.
-We find that planning is disincentivized later, but only for easier levels.
-We offer a hypothesis about why the training process disincentivizes the planning effect: the NN finds levels easier (needs less thinking), and also learns when to do the thinking it needs (via pacing).
-7 Acknowledgements
-We would like to thank ChengCheng Tan for editing this paper, Philip Quirke for project management, and the rest of the FAR team for general support.
-We also thank Maximilian Li, Alex F.
-Spies, and anonymous reviewers for improving the paper with their comments and edits.
-We are also very grateful for the many questions about their RL setup that Arthur Guez and Timothy Lillicrap answered, as well as inspiration from Stephen Chung and Scott Goodfriend.
-Appendix A Training hyperparameters
+A.1 Training hyperparameters
 All networks were trained with the same hyperparameters, which were tuned on a combination of the ResNet and the $\text{DRC}(3,3)$.
 These are almost exactly the same as Guez et al.
-(2019), allowing for taking the mean of the per-step loss instead of the sum.
+[2019], allowing for taking the mean of the per-step loss instead of the sum.
+Time limits.
+During training, we want to prevent strong time correlations between the returns, so the gradient steps are not correlated over time.
+For this reason, the time limit for each episode is uniformly random between 91 and 120 time steps.
 Loss.
@@
 The original IMPALA implementation, as well as Huang et al.
-(2023), sum the per-step losses.
+[2023], sum the per-step losses.
 We instead average them for more predictability across batch sizes, so we had to scale down some parameters by a factor of $1/640$: Adam $\epsilon$, gradient norm for clipping, and L2 regularization).
 Weight initialization.
-We initialize the network with the Flax (Heek et al., 2023) default: normal weights truncated at 2 standard deviations and scaled to have standard deviation $\sqrt{1/\text{fan\_in}}$.
+We initialize the network with the Flax [Heek et al., 2023] default: normal weights truncated at 2 standard deviations and scaled to have standard deviation $\sqrt{1/\text{fan\_in}}$.
 Biases are initialized to 0.
-The forget gate of LSTMs has 1 added to it (Jozefowicz et al., 2015).
+The forget gate of LSTMs has 1 added to it [Jozefowicz et al., 2015].
 We initialize the value and policy head weights with orthogonal vectors of norm 1.
@@
 We believe this has essentially no effect, but we left it in to more closely match Guez et al.
-(2019).
+[2019].
 Software.
-We base our IMPALA implementation on Cleanba (Huang et al., 2023). We implemented Sokoban in C++ using Envpool (Weng et al., 2022) for faster training, based on gym-sokoban (Schrader, 2018).
-A.1 Number of training steps
+We base our IMPALA implementation on Cleanba [Huang et al., 2023]. We implemented Sokoban in C++ using Envpool [Weng et al., 2022] for faster training, based on gym-sokoban [Schrader, 2018].
+A.2 Number of training steps
 […]
@@
 […]
-Appendix B Learning curve comparison
+A.3 Learning curve comparison
 It is difficult to fully replicate the results by Guez et al.
-(2019).
+[2019].
 Chung et al.
-(2024) propose an improved method for RL in planning-heavy domains.
+[2024] propose an improved method for RL in planning-heavy domains.
 They employ the IMPALA $\text{DRC}(3,3)$ as a baseline and plot its performance in Chung et al.
-(2024, Figure 5).
+[2024, Figure 5].
 They plot two separate curves for $\text{DRC}(3,3)$: that from Guez et al.
-(2019), and a decent replicated baseline.
+[2019], and a decent replicated baseline.
 The baseline is considerably slower to learn and peaks at lower performance.
@@
 We compare our replication to Guez et al.
-(2019) in Figure 8, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1).
+[2019] in Figure 7, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1).
 Our implementation also appears much less stable, with large error bars and large oscillations over time.
@@
 The parameter counts (Table 4) are very different from what Guez et al.
-(2019) report.
+[2019] report.
 In private communication with the authors, we confirmed that our architecture has a comparable number of parameters, and some of the originally reported numbers are a typographical error.
@@
 $\text{DRC}(3,3)$ 1,285,125 (1.29M)
-DRC(1,1) 987,525 (0.99M)
+$\text{DRC}(1,1)$ 987,525 (0.99M)
 ResNet: 3,068,421 (3.07M)
@@
 Each architecture has 5 random seeds, the solid line is the pointwise median and the shaded area spans from the minimum to the maximum.
-The dotted lines are data for the performance of architectures extracted from the (Guez et al., 2019) PDF file.
-The values are slightly different from what Figures 1 and 3 report because they are calculated on a random sample of 1024 levels (24 levels are repeated for test-unfiltered).
+The dotted lines are data for the performance of architectures extracted from the [Guez et al., 2019] PDF file.
+The values are slightly different from what Figures 5 and 3 report because they are calculated on a random sample of 1024 levels (24 levels are repeated for test-unfiltered, which only has 1000 levels).
 Table 5: Success rate and return of DRC and ResNet on the unfiltered test set at various training environment steps.
@@
 2B 97.9 10.29 99.3 10.52 59.4 1.16 76.6 4.52
-Appendix C Additional cycles investigation
-We run the DRC(3,3) on medium-validation levels and record where cycles in the game state happen.
-We prune out the redundant cycles that visit the same states as a given cycle.
-There are in total 13702 non-redundant cycles.
-For each cycle, if it is of length $N$, we replace it with $N$ steps of thinking, and measure what happens.
-In 86.74% of cases the DRC(3,3) does not immediately start going in cycles after the thinking time.
-That is, no cycle begins in the step after thinking time.
-In 82.39% of cases this persists for the next $N$ time steps: no cycles start in the $N$ steps after thinking.
+Appendix B Generalizing the $\text{DRC}(3,3)$ to larger levels
+We license the levels by Þorsteinsson [2009] as GPLv3, and make them available in the supplementary material.
+The performance of the $\text{DRC}(3,3)$ on each set of levels is in Table 6.
+The sets with higher scores are also those which humans find easier: for example, “Dimitri & Yorick” was made for children by Jacques Duthen and consists entirely of relatively small levels (largest is $12\times 10$) with at most 5 boxes.
+The sets where the $\text{DRC}(3,3)$ solves nothing are also difficult for the authors of this paper.
+We tried giving the $\text{DRC}(3,3)$ from 2 to 128 extra thinking steps for these levels, incrementing in powers of two.
+For most sets, we find some benefit to 2-4 thinking steps, but no more.
+The sole exception is XSokoban, which contains one level (Figure 2, right) which requires 128 steps of thinking to solve.
+We encourage the reader to go to the website by Þorsteinsson [2009] and try solving the levels.
+(a) Microban, level 105
+(b) Microban, level 144
+(c) Mas Sasquatch, level 15
+Figure 8: Thinking steps are not always useful: For level (a), the network succeeds with 0 to 16 thinking steps, but fails with 32 or more thinking steps.
+Levels (b) and (c) are too difficult for the $\text{DRC}(3,3)$, and it always fails on them independent of thinking steps, though it comes up with decent partial solutions.
+All levels from Þorsteinsson [2009].
+(a) Sokoban Jr. 2, level 17
+(b) Microban, level 145
+(c) XSokoban, level 30
+Figure 9: Some levels require thinking steps for success: For level (a), the network always succeeds.
+For level (b), it succeeds at 32 or more thinking steps.
+For level (c), it needs 128 thinking steps to succeed.
+All levels from Þorsteinsson [2009].
+Table 6: Performance of the $\text{DRC}(3,3)$ on each set of levels by [Þorsteinsson, 2009].
+The “Max solved” columns represent the proportion of levels solved at the number of steps in the “at steps” column, which is the highest solved proportion for each number of thinking steps tried.
+All levels Levels larger than $10\times 10$
+Level collection # solved Max slv. max at # solved Max slv. max at
+Dimitri & Yorick 61 83.6% 88.5% 16 0
+Sokoban Jr. 1 60 80.0% 81.7% 2 19 73.7% 73.7% 0
+Howard’s 3rd set 40 60.0% 62.5% 4 1 0.0% 0.0% 0
+Simple sokoban 61 54.1% 60.7% 16 51 47.1% 54.9% 16
+Sokoban Jr. 2 54 48.1% 50.0% 2 40 45.0% 47.5% 2
+Microban 155 23.9% 24.5% 4 17 5.9% 11.8% 2
+Deluxe 55 21.8% 23.6% 2 1 0.0% 0.0% 0
+Sokogen 990602 78 20.5% 23.1% 8 0
+Yoshio Automatic 52 13.5% 17.3% 2 0
+Sasquatch III 16 6.2% 6.2% 0 8 0.0% 0.0% 0
+Howard’s 1st set 100 5.0% 6.0% 2 54 0.0% 0.0% 0
+Howard’s 2nd set 40 5.0% 10.0% 4 22 0.0% 0.0% 0
+Microcosmos 40 5.0% 7.5% 4 0
+Still more levels 35 2.9% 2.9% 0 34 2.9% 2.9% 0
+Sasquatch IV 36 2.8% 2.8% 0 20 0.0% 0.0% 0
+Xsokoban 40 2.5% 5.0% 128 39 2.6% 5.1% 128
+Sasquatch 49 2.0% 4.1% 4 39 0.0% 2.6% 4
+David Holland 1 10 0.0% 0.0% 0 5 0.0% 0.0% 0
+David Holland 2 10 0.0% 0.0% 0 9 0.0% 0.0% 0
+Howard’s 4th set 32 0.0% 0.0% 0 30 0.0% 0.0% 0
+Mas Sasquatch 50 0.0% 0.0% 0 43 0.0% 0.0% 0
+Nabokosmos 40 0.0% 0.0% 0 0
+Sokoban 50 0.0% 0.0% 0 48 0.0% 0.0% 0
+Appendix C Case Studies
+(a) file 0, level 18
+[…]
+(b) file 0, level 53
+[…]
+(c) file 0, level 153
+[…]
+Figure 10: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps.
+Colors are as in Figure 12.
+Boxes and targets are paired in upper- and lower-case letters respectively, such that the DRC’s best solution places boxes on targets in alphabetical order.
+Videos available at this https URL.
+Levels solved faster obtain higher return as the agent incurs the per-step penalty fewer times.
+Note that the letters used are for reference and not a property of Sokoban.
+Case Study: Thinking makes some levels solvable Figure 10(a).
+For example, thinking lets the DRC solve Figure 10(a).
+In the no-thinking condition, the DRC first pushes box $C$ one square to the right.
+It then goes back to push $A$ to $a$, but it has now become impossible to push box $B$ onto $b$.
+In contrast, after thinking, the DRC pushes $A$ to $a$ first, which lets it solve the level.
+Case Study: Thinking speeds up solving Figure 10(b).
+Figure 10(b) illustrates a scenario where the agent happens.
+In the no-thinking condition, the DRC takes many steps back and forth before pushing any boxes.
+First it goes down to $y$, up to $c$, then down onto $z$, back up to $y$ and to $z$ again.
+It then proceeds to solve the rest of the puzzle: push box $A$ onto $a$, prepare box $B$ on $x$ and box $C$ where $B$ originally was, push in boxes $B,C$ and finally $D$.
+In the thinking condition, the DRC makes a beeline for $A$ and then plays the same solution.
+Case Study: Thinking slows downFigure 10(c).
+Figure 10(c) illustrates a scenario where thinking time results in a slower solution.
+In the no thinking condition, the DRC starts by pushing box C into position y, then pushes boxes A, then B into place.
+On the way back down, the DRC pushes C onto c and finally D onto d.
+In contrast, in the thinking condition, the DRC goes the other way and starts by pushing B onto b.
+The subsequent solution (A, C, then D) is the same, but the DRC has wasted time trekking back from B to A.
+Appendix D Probe Training
+Probe architecture
+We train multiple different linear probes on the hidden states $h$ and cell states $c$ activations of the network.
+The probes are trained either on the states of the individual layers or by concatenating the states of all the three layers.
+We collect the activations of the model by letting it play through the hard levels.
+The training set for the probes are constructed by taking a random sample of cached levels and including all the timesteps in the level excluding the first five timesteps.
+We exclude these intial steps as we believe they can be noisy as the network comes up with the stategy to solve the levels.
+The probes are trained with logistic regression with L1 decay using the Scikit-Learn library.
+We do a grid-search on the learning rate and the L1 weight decay to select the probe that has the highest F1-score on a validation set.
+For multi-class probe targets, each potential output class $l$ is treated like a different data point.
+That is: for each label $l$, we consider a prediction positive if the highest probe logit is for label $l$, and negative otherwise.
+We consider a data point positive iff the true label is label $l$.
+Then, we compute the confusion matrix and the F1 score from the $n\cdot l$ data points.
+Appendix E Details about Sparse Autoencoders (SAEs)
+Figure 11: Visualization of some interpretable features from the SAE of last layer.
+These features also appear monosemantically in the channels.
+The precision, recall, and F1 score for the features are reported in Table 10.
+In order to search for more monosemantic and interpretable features in the network, we train sparse-autoencoders (SAEs) [Huben et al., 2023, Bricken et al., 2023] on the individual squares in the $h$ hidden state of the network consisting of 32 neurons.
+Thus, we get a $10\times 10$ visualization for each SAE feature as shown in Figure 11.
+We use the top-k activation function [Gao et al., 2024, Tamkin et al., 2024] for the SAE as it is currently the state-of-the-art method to train SAEs and is easier to work with as it directly sets the $L_{0}$ norm of the SAE activations.
+The hyperparameter search space for training the SAE is provided in Table 7.
+We train separate SAEs for each layer with the specified hyperparameters and pick the one that achieves greater than $90\%$ explained variance while having interpretable features assessed through manual visual inspection.
+We release these trained SAEs and probes in the same huggingface repo as our trained DRC networks.
+[^8] https://huggingface.co/AlignmentResearch/learned-planner
+We provide examples of some of the interpretable features from an SAE trained on the last layer with $k=8$ in Table 8 with corresponding visualization in Figure 11.
+For the “Target”, “Unsolved”, and “Solved” concepts (Table 10), we observed cases where the SAE feature is offset from the ground truth by 1 square in either horizontal or vertical directions.
+The level setup includes a permanent one-square outer-edge wall, so this offset never results in an out-of-bounds issue.
+We evaluated these potential “Offset” variants for the Target, Unsolved, and Solved concepts.
+All of the features in the SAE that we find to be interpretable are already embedded in individual channels.
+These channels are either as monosemantic as the SAE features, or in some cases more monosemantic as measured by the F1 score.
+Table 3 reports the precision, recall, and F1 scores for the action features as evaluated through channels, SAE neurons, and linear probe trained against the ground truth action predictions.
+Table 10 reports the same scores for additional interpretable features.
+The SAE action features have lower F1 scores than channels by a margin of $5.9\%$ on average.
+On the other hand, the linear probes trained across all channels of the hidden state have similar F1 scores as the specified channels indicating that the channels are monosemantic and cannot be improved upon by combining various channels using a linear probe.
+Table 7: Hyperparameter search space for training SAE
+Hyperparameter Search Space
+$k$$\{4,8,12,16\}$
+learning rate $\{1e-5,5e-5,1e-4,5e-4,1e-3\}$
+expansion factor $\{16,32,64\}$
+Table 8: SAE Feature Concepts
+Concept Description
+Target The 4 target squares (static)
+Unsolved Targets and boxes that aren’t solved
+Solved Solved target squares with a box on them
+Agent Up The agent will move Up next step
+Agent Down The agent will move Down next step
+Agent Left The agent will move Left next step
+Agent Right The agent will move Right next step
+Table 9: Breakdown of levels by category at 6 thinking steps.
+Level categorization Percentage
+Solved, previously unsolved 6.87
+Unsolved, previously solved 2.23
+Solved, with better returns 18.98
+Solved, with the same returns 50.16
+Solved, with worse returns 5.26
+Unsolved, with same or better returns 15.14
+Unsolved, with worse returns 1.36
+Table 10: Scores for SAE and Channel features
+Concept Offset $(dy,dx)$ Channel SAE Feature
+Number Prec Rec F1 Number Prec Rec F1
+Target (1, 0) L3C17 97.8 97.7 97.8 L3F278 97.8 98.1 98.0
+Unsolved targets and boxes (0, 0) L3C7 94.9 90.8 92.8 L3F212 95.3 86.6 90.7
+Solved targets (0, 0) -L3C7 91.6 94.6 93.0 L3F179 91.7 91.5 91.6
+Appendix F Additional quantitative behavior figures and tables
+Figure 12: Left: A Sokoban level from the hard set. Right: the same level as the NN sees it, one pixel per tile. Walls are black, boxes are brown, targets are pink and the robot is green.
+Figure 13: Number of thinking steps required to solve the level vs.
+number of nodes A* needs to expand to solve it.
+The trend is increasing at the end but very unclear, indicating different heuristics used by the DRC and A*.
+Figure 14: Success rate on datasets of various difficulty, for various checkpoints of each architecture.
+This deviates very little from a curve, which shows that ResNets and DRCs which are equally good at the easier sets are also equally good at the harder sets.
+Perhaps DRC(1,1) is a slight exception, but it also performs much worse than the others overall (see Figure 7).
 (a) Cycle length distribution
@@
 (c) with cycles from those levels that were solved by the agent
-Figure 9: We replace N-length cycles with N thinking steps and checking for the same state after some timesteps.
+Figure 15: We replace $N$-length cycles with $N$ thinking steps and check for the same state after some timesteps.
 (a) A histogram of cycle lengths in the medium-validation set.
 (b, c) After replacing a cycle with the same length in thinking steps, are all the states the same for the next $x$ steps?
-Appendix D Additional quantitative behavior figures
-Figure 10: Success rate on datasets of various difficulty, for various checkpoints of each architecture.
-This deviates very little from a curve, which shows that ResNets and DRCs which are equally good at the easier sets are also equally good at the harder sets.
-Perhaps DRC(1,1) is a slight exception, but it also performs much worse than the others overall (see Figure 8).
-Figure 11: Number of thinking steps required to solve the level vs.
-number of nodes A* needs to expand to solve it.
-The trend is somewhat increasing but much less clear, indicating different heuristics used by the NN and A*.

### v3

@@
 Abstract
-How a neural network (NN) generalizes to novel situations depends on whether it has learned to select actions heuristically or via a planning process.
+Planning is essential for solving complex tasks, yet the internal mechanisms underlying planning in neural networks remain poorly understood.
+Building on prior work, we analyze a recurrent neural network (RNN) trained on Sokoban, a challenging puzzle requiring sequential, irreversible decisions.
+We find that the RNN has a causal plan representation which predicts its future actions about 50 steps in advance.
+The quality and length of the represented plan increases over the first few steps.
+We uncover a surprising behavior: the RNN “paces” in cycles to give itself extra computation at the start of a level, and show that this behavior is incentivized by training.
+Leveraging these insights, we extend the trained RNN to significantly larger, out-of-distribution Sokoban puzzles, demonstrating robust representations beyond the training regime.
+We open-source our model and code, and believe the neural network’s interesting behavior makes it an excellent model organism to deepen our understanding of learned planning.
+[…]
+1 Introduction
+In many tasks, the performance of both humans and some neural networks (NNs) improves with more reasoning: whether by giving a human time to think before making a chess move, or by prompting or training a large language model (LLM) to reason step by step [Kojima et al., 2022, OpenAI, 2024].
+Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment.
+So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning [Hubinger et al., 2019] – may internalize goals different from the training objective, leading to goal misgeneralization [Di Langosco et al., 2022, Shah et al., 2022].
+Understanding how NNs learn to plan and represent the objective could be key to detect, prevent or correct goal misgeneralization.
+In this work, we focus on interpreting a Deep Repeating ConvLSTM [Guez et al., 2019, DRC] trained on Sokoban, a puzzle game often used as a planning benchmark [Peters et al., 2023].
+We interpret the best network from Guez et al.
+[2019], $\text{DRC}(3,3)$, with 3 recurrent layers that are applied 3 times per environment step.
+Further details of the network are provided in Section 2.
+We find that its internal plan representation [Bush et al., 2025] is causal, improves with more computation, and that the DRC learns to take advantage of that by often “pacing” to get enough time to refine its internal plan.
+We show similar results in Appendix B for another DRC network and causal plan representation in a ResNet model.
+1.1 Definitions
 Guez et al.
-[2019, “An investigation of model-free planning”] found that a recurrent NN (RNN) trained to play Sokoban appears to plan, with extra computation steps improving the RNN’s success rate.
-We replicate and expand on their behavioral analysis, finding the RNN learns to give itself extra computation steps in complex situations by “pacing” in cycles.
-Moreover, we train linear probes that predict the future actions taken by the network and find that intervening on the hidden state using these probes controls the agent’s subsequent actions.
-Leveraging these insights, we perform model surgery, enabling the convolutional NN to generalize beyond its $10\times 10$ architectural limit to arbitrarily sized inputs.
-The resulting model solves challenging, highly off-distribution levels.
-We open-source our model and code, and believe the neural network’s small size (1.29M parameters) makes it an excellent model organism to deepen our understanding of learned planning.
-[…]
-Figure 1: Top left: Proportion of medium-difficulty validation levels solved vs.
-environment steps used in training.
-Curves show the RNN’s performance with a specific number of thinking steps (forced no-op actions) at episode start, along with a ResNet baseline.
-Bottom left: Estimated planning effect: the 8-steps minus the 0-steps curve.
-We see that reasoning emerges in the first 70M steps and keeps increasing for the hard levels (red), but decreases for medium levels (blue).
-Right: Linear probe predictions (arrows) for the direction the box will move in.
-Opacity is proportional to the number of steps in the episode in which the probe predicted a particular arrow.
-Correct predictions are in green and incorrect in red, the red predictions show other plausible plans for Box 4.
-This probe causally affects the actions of the agent, as described in Section 4.
-Boxes are numbered in the order in which they arrive at their final target.
-1 Introduction
-In many tasks, the performance of both humans and sufficiently advanced neural networks (NNs) improves with more reasoning: whether by giving a human more time to think before taking a chess move, or by prompting or training a large language model (LLM) to reason step by step [Kojima et al., 2022, OpenAI, 2024].
-Among other reasoning capabilities, goal-oriented reasoning is particularly relevant to AI alignment.
-So-called “mesa-optimizers” – AIs that have learned to pursue goals through internal reasoning [Hubinger et al., 2019] – may internalize goals different from the training objective, leading to harmful misgeneralization [Di Langosco et al., 2022, Shah et al., 2022].
-Within goal-oriented reasoning, this work distinguishes between plans and search algorithms.
-A plan is a sequence of (perhaps abstract) actions which an agent intends to take.
-A search algorithm is an algorithm that considers many plans, evaluates them according to their predicted outcomes, and picks the plan evaluated as best.
-Search algorithms perform better the more plans they can evaluate and, with enough compute, generalize very well to novel problem instances [Russell and Norvig, 2009].
-This work investigates a recurrent neural network (RNN) that plays Sokoban [Guez et al., 2019], a challenging puzzle game that remains a benchmark for planning algorithms [Peters et al., 2023] and reinforcement learning [Chung et al., 2024].
-Using linear location-wise probes from concurrent work [Bush et al., 2024], we show that its activations represent a plan in a way which causes the RNN’s actions.
-The behavior of the RNN suggests that it also performs search.
-We find that the RNN sometimes takes actions which return to previous states of the environment.
-This overwhelmingly happens at the beginning of episodes and can be substituted by no-ops, suggesting the RNN deliberately goes in cycles to get more time to think.
-Giving the RNN time to think with no-ops [Guez et al., 2019] disproportionately helps with levels that are difficult (Figure 1), long, or that require waiting many steps before the first reward (Figure 3).
-Using insights from the plan probes and inspecting activations, we perform model surgery to uncover latent generalization capabilities of the RNN.
-The RNN is trained only on $10\times 10$ levels with 4 boxes (and its output MLP can only handle $10\times 10$ levels), but we are able to make it play on larger, nontrivial levels with more boxes (Figure 2).
-Altogether, evidence from probes and behavior show that the RNN causally represents a plan.
-The evidence highly suggests that the RNN implements an online search algorithm which scales with computation, although we cannot rule out multi-step heuristics.
-Consequently, we believe this RNN is an excellent model organism for understanding planning in sophisticated neural networks.
-Its behavior is complex enough to be interesting and its size (1.29M parameters) is small enough to be tractable to reverse engineer.
-This work shows that supervised probes and sparse autoencoders [Huben et al., 2023, Bricken et al., 2023, Gao et al., 2024] are not enough to fully do so, but it is likely that other mechanistic interpretability techniques such as circuit analysis [Olah et al., 2020, Geiger et al., 2021, Conmy et al., 2023] will reveal more about the algorithm which produces plans.
-Figure 2: The $\text{DRC}(3,3)$ is limited to $10\times 10$ grids by a fixed-dimension MLP.
-With model surgery, the ConvLSTM backbone generalizes to much more challenging settings than the $10\times 10$ training set grids.
-Left: XSokoban-31, which the $\text{DRC}(3,3)$ solves with 0 thinking steps.
-Right: XSokoban-29, which the $\text{DRC}(3,3)$ solves but only after 128 thinking steps.
-More examples in Appendix B.
+[2019] showed that the DRC is very capable, generalizes well, and its performance improves with thinking steps at the beginning of an episode.
+Based on this, they claimed that DRC internally plans, but did not make precise what this means.
+To clarify our contributions, we introduce a distinction between plans and search algorithms.
+Definition 1.1 (Plan).
+A plan is a sequence of future actions $\{a_{t}\text{ for all }t>t_{0}\}$.
+Definition 1.2 (Causally represented plan).
+Let $m\in\cal{M}$ be a simple model (e.g. logistic regression). A plan is causally represented if, for the current neural state $z_{t}$, and a hypothetical future environment state $s_{k}$, $k>t$:
+  1. 1.
+    Prediction: we can extract the plan using $m$ with sufficient accuracy: $a_{k}\approx m(z_{t},s_{k})$ on average for $k>t$.
+  2. 2.
+    Causality: modifying $z_{t}$ alters the plan according to $m$, that is, preserves $a_{k}\approx m(z_{t},s_{k})$.
+Definition 1.3 (Search).
+A search algorithm is a decision-making process that involves generating multiple possible partial or complete plans, evaluating their predicted outcomes, and selecting the plan with the highest value.
+Unlike traditional planning algorithms, search in neural networks may use heuristics or partial models to evaluate plans, rather than specifying an exact world state they correspond to.
+The key feature of search is the explicit representation and selection among competing plans.
+The competing plans do not have to be complete and could be a partial rollout.
+Definition 1.4 (Thinking steps).
+Timesteps where the agent receives the same observation repeatedly, with its predicted actions not executed in the environment.
+Definition 1.5 (Cycle).
+A sequence of environment steps that starts and ends at the same state.
+1.2 Contribution statement
+Finding 1 (The DRC causally represents its plan).
+The DRC maintains a plan that is consistently selected and causally represented in its hidden states.
+This plan predicts and drives the agent’s behavior, and intervening on these representations alters the agent’s actions.
+See Section 3.
+Concurrently, Bush et al.
+[2025] also demonstrated finding 1 on handcrafted toy levels similar to the ones in Appendix D, by intervening on a key step.
+We further test the causality of similar representations on random perturbations on difficult levels and propose a more precise causal intervention method, which works on any level and lets us alter any step in a plan.
+Finding 2 (Plan improves with computation).
+With each internal iteration, the DRC network refines its plan, akin to doing a heuristic search, bringing it closer to the agent’s eventual sequence of actions. See Section 4.
+Finding 3 (Pacing behavior).
+The DRC sometimes delays irreversible actions to allocate more computation time for refining its plan. This behavior is discovered during training and is not an artifact. See Section 5.
+Evidence for 3: the DRC’s probed plans change 60% faster during cycles compared to non-cycle steps.
+Additionally, training the DRC with a bonus for NOOPs (smaller than the per-step penalty) increases the proportion of NOOPs in cycle steps while keeping the total number of cycle steps the same.
+We believe finding 3 is likely the correct explanation for Guez et al.
+[2019]’s finding that forced thinking time improves performance.
+Finding 4 (Generalization to OOD larger levels.).
+Using action probes, we are able to make the convolutional core of the DRC generalize to novel, significantly larger Sokoban puzzles beyond its training distribution. See Section 6.
+Previous work takes significant out-of-distribution generalization as evidence of algorithmic reasoning [Bansal et al., 2022, Guez et al., 2019].
+Finding 5 (Lack of explicit search, low confidence).
+The $\text{DRC}(3,3)$ does not perform search by generating and evaluating multiple plans, before selecting the one with the highest predicted value.
+We were unsuccessful at finding a representation of the value of each plan using a value probe (Appendix F), and at correlating thinking steps with A* expanded nodes (Figure 21).
+It is possible that DRC performs search but with a different heuristic.
+The rest of the evidence is compatible with search, and with the DRC iteratively refining its plan with heuristics.
+We demonstrate these findings on $\text{DRC}(3,3)$ in the main text.
+In Appendix B, we show that the major Findings 3--3 also hold true for $\text{DRC}(1,1)$, while only Finding 3 holds true for a ResNet model.
+Open-source resources.
+We open-source all models, tools, and interpretability data, offering a comprehensive resource for future research into planning behaviors [^1] URL references removed during double-blind review.
+Code is available in the supplementary material..
+The DRC achieves an ideal balance of complexity and tractability for interpretability research with just 1.29M parameters.
+Refer to caption  Refer to caption
+Figure 1: Top left: Validation success rate on medium-difficulty levels solved vs.
+training steps, with varying numbers of initial thinking steps (forced NOOP actions) and a ResNet baseline.
+Bottom left: Estimated planning effect (8-steps minus 0-steps) showing that planning emerges within 70M steps and increases for the hard levels (red) but decreases for medium levels (blue).
+Middle: Linear probe predictions for box movement direction.
+Green arrows indicate correct predictions and incorrect red arrows indicate plausible alternatives for Box 4.
+Opacity reflects frequency during the episode.
+This probe causally affects agent actions as described in Section 3.
+Boxes are numbered in their order of placement on targets.
+Right: The same level’s input representation to the NN, where each pixel is a single tile.
+Walls are black, boxes are brown, targets are pink, and the robot is green.
+Figure 2: Left: The $\text{DRC}(3,3)$ architecture [Guez et al., 2019] with 3 convolutional layers repeated 3 times, embedded observation feeding into each layer, the last layer’s output $h$ feeding back into the first layer, and recurrent connections between layers.
+Blue: Linear probes on hidden state grid cells to predict future actions (Section 3).
+Red: Replacing the fixed-dimension MLP with probes enables the ConvLSTM core to generalize to more challenging puzzles beyond the $10\times 10$ training grids (Section 6).
+Right: $13\times 17$ level from XSokoban-31 is solved by $\text{DRC}(3,3)$ after replacing the MLP.
 2 Setting up the test subject
 We train an agent closely following the setup from Guez et al.
-[2019], using the IMPALA [Espeholt et al., 2018] reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture.
+[2019], using the IMPALA V-trace actor-critic [Espeholt et al., 2018] reinforcement learning (RL) algorithm with Guez et al.’s Deep Repeating ConvLSTM (DRC) recurrent architecture.
 We also train a ResNet baseline.
-We open-source both the trained networks  [^1] https://github.com/AlignmentResearch/learned-planner and training code  [^2] https://github.com/AlignmentResearch/train-learned-planner .
-For more details about architectures and training, see Appendix A.
+For further architectural and training details, see Appendix A.
 $\text{DRC}(D,N)$ architecture.
-Most of this paper focuses on the behavior and representations of a $\text{DRC}(3,3)$ neural network [Guez et al., 2019].
-The core of this network is a $D$-layer ConvLSTM [Shi et al., 2015], which is repeatedly applied $N$ times per environment step.
-The output of the last ConvLSTM layer (the $D$th layer) is fed to the input at the next tick, effectively giving the network $D\cdot N$ layers of sequential computation to output the next action.
-We take $D=3,N=3$.
-A linear combination of the mean- and max-pooled activations of the ConvLSTM is also fed to the next step, thus letting it communicate quickly across the receptive field (known as pool-and-inject).
-An encoder block consisting of two $4\times 4$ convolutions processes the input, which is also fed to each ConvLSTM layer.
-At the end, an MLP with 256 hidden units transforms the flattened ConvLSTM outputs into the policy (actor) and value function (critic) heads.
+This paper primarily focuses on the behavior and representations of a $\text{DRC}(3,3)$ neural network [Guez et al., 2019].
+The core component of this network is a $D$-layer ConvLSTM [Shi et al., 2015], which is repeatedly applied $N$ times per environment step (Figure 2, left).
+The output of the final ConvLSTM layer (the $D$th layer) is fed back into the input of the first layer at the next tick, effectively giving the network $D\cdot N$ layers of sequential computation to determine the next action.
+In our setup, $D=N=3$ and $C=32$.
+Appendix B contains results on $\text{DRC}(1,1)$ and a ResNet model.
+A linear combination of the mean- and max-pooled ConvLSTM activations is injected into the next step, enabling quick communication across the receptive field, known as (pool-and-inject).
+An encoder block consisting of two $4\times 4$ convolutions process the input, which is fed to each ConvLSTM layer.
+Each ConvLSTM layer’s hidden states $(h,c)$ have the same number of channels $C$.
+Finally, an MLP with 256 hidden units transforms the flattened ConvLSTM outputs into the policy (actor) and value function (critic) heads.
 Dataset.
-Sokoban is a grid puzzle game with walls, floors, movable boxes, and target tiles.
-The player’s goal is to push all boxes onto target tiles while navigating walls.
-We use the Boxoban dataset [Guez et al., 2018], consisting of $10\times 10$ procedurally generated Sokoban levels, each with 4 boxes and targets.
+Sokoban is a grid-based puzzle game with walls, floors, movable boxes, and target tiles.
+The goal is to push all boxes onto target tiles while avoiding obstacles.
+We use the Boxoban dataset [Guez et al., 2018], consisting of $10\times 10$ procedurally generated levels, each with 4 boxes and targets.
 The edge tiles are always walls, so the playable area is $8\times 8$.
-Boxoban separates levels into train, validation and test sets, with three difficulty levels: unfiltered, medium, and hard.
+Boxoban separates levels into train, validation, and test sets with three difficulty levels: unfiltered, medium, and hard.
 Guez et al.
-[2019] generated these sets by filtering levels that cannot be solved by progressively better-trained DRC networks, so easier sets occasionally contain difficult levels.
-In this paper, we use the unfiltered-train (900k levels) set to train networks.
-To evaluate them, we use the unfiltered-test (1k levels) [^3] We use unfiltered-test rather than unfiltered-validation so the numbers are directly comparable to Guez et al.
+[2019] generated these sets by filtering levels unsolvable by progressively better-trained DRC networks.
+So easier sets occasionally contain difficult levels.
+In this paper, we train agents on the unfiltered-train (900k levels).
+For evaluation, we use the unfiltered-test (1k levels) [^2] We use unfiltered-test rather than unfiltered-validation to ensure direct comparability with Guez et al.
 [2019]., medium-validation (50k levels), and hard ($\sim$3.4k levels) sets, which do not overlap.
-To test $\text{DRC}(3,3)$ generalization to different sizes, we use the levels collected by Þorsteinsson [2009] (see Appendix B).
+To test $\text{DRC}(3,3)$ generalization to different sizes, we use the levels collected by Þorsteinsson [2009] (see Appendix C).
 Environment.
-The observations are $10\times 10$ RGB images, normalized by dividing each component by 255.
-Each type of tile is represented by a pixel of a different color [Schrader, 2018], an example is in Figure 12.
-The player can move in cardinal directions (Up, Down, Left, Right).
-The reward is -0.1 per step, 1 for placing a box on a target, -1 for taking it off, and 10 for finishing the level by placing all of the boxes.
-The time limit for evaluation is 120 steps, though large levels in Section 5.1 use 1000 steps.
-3 Behavioral evidence of search
-3.1 Effects of thinking time: non-myopic network
-First, we examine the effect of thinking time: introducing no-ops at the beginning of an episode, while letting the RNN process its hidden state.
-As in Guez et al.
-[2019], adding 6 extra thinking steps lets the $\text{DRC}(3,3)$ solve an extra $4.7\%$ of levels, slightly decreasing when going up to 16 steps (Figure 5, left).
-We find that the extra thinking time disproportionately helps with more difficult levels, as measured by longer optimal solutions (Figure 3, right).
-We also checked whether thinking time correlates with the number of nodes expanded by an A* search, but it does not (Figure 14).
-Thinking time helps varying amounts depending on training time and level difficulties: Figure 1 (bottom left) shows that most of the planning effect steeply appears during the first 70M steps of learning.
-Over the remaining training time, the RNN benefits more from thinking steps for the hard levels, but less for the medium levels; indicating perhaps that it learns better heuristics for them.
-We also find that thinking time disproportionately helps with levels in which the agent can greedily put a box to target early, but where doing so is detrimental to solving the level.
-Figure 3 (left, b) shows this: with 0 thinking steps, the first box is pushed to target on average 4 steps too early; compared to the solution reached at 6 thinking steps.
-Some of this is driven by the forced thinking steps preventing the $\text{DRC}(3,3)$ from taking a catastrophic action in the first few moves.
-However, the average time to first box is more than 6, even for 0 thinking steps, so there must also be many cases in which the 6 extra thinking steps give the NN time to find a better solution, rather than just preventing it from greedily pushing the first box.
-It must also be less greedy for boxes 2-3.
+The observations are $10\times 10$ RGB images, normalized by dividing each pixel component by 255.
+Each tile type is represented by a unique pixel color [Schrader, 2018], illustrated in Figure 1 (right).
+The player has four actions available to move in cardinal directions (Up, Down, Left, Right).
+The reward is -0.1 per step, +1 for placing a box on a target, -1 for removing it, and +10 for finishing the level by placing all of the boxes.
+The time limit for evaluation is 120 steps, although large levels in Section 6 use 1000 steps.
+In the NOOP training part of Section 5, we modify the environment by adding an explicit NOOP action that has a fraction of the penalty as the other action.
+3 The $\text{DRC}(3,3)$ causally represents its plan
+We train probes [^3] In the interpretability literature, probes are simple (usually linear) models that are trained to predict specific labels (referred to as ‘concepts’) from intermediate activations of a NN [Alain and Bengio, 2016, Belinkov, 2016].
+They provide a way to decode information represented in the network.
+However, that a probe can predict some information from NN activations is not enough to show that the NN is using that information.
+To show that, Li et al.
+[2023] argue that one has to intervene on the NN activations to change the probe output, and observe the NN’s behavior changing accordingly.
+to predict the future actions of $\text{DRC}(3,3)$ and other features of the environment.
+Of these, the probe predicting the future move-direction of boxes from every square in the grid has the strongest causal effect on the actions of the DRC: intervening on the activations to change the probe’s output also affects the future box directions [Li et al., 2023].
+Qualitatively, the sequence of box movements contains most of the information needed to solve Sokoban levels, so we consider this a plan.
+Thus, we consider this strong evidence that the $\text{DRC}(3,3)$ represents and uses plans (finding 1).
+The spatial structure of the probes and some of the targets are adapted from Bush et al.
+[2025].
+We find some evidence that the $\text{DRC}(3,3)$ considers multiple plans.
+For example, in the level in Figure 1 (middle), the box probe initially predicts moving box 4 down and right (shown in red) but later revises this plan and takes different actions.
+Despite this, we have not found evidence of multiple simultaneous plan representations or a mechanism to evaluate and decide between plans, such as a value comparison.
+Overall, this section provides strong evidence for a causally represented plan (finding 1), but very limited evidence supporting search (finding 5).
+3.1 Probe methodology: interpretation and intervention
+Training.
+We train linear regression probes with L1 decay to predict environment features from the agent’s activations.
+Similar to the concurrent work by Bush et al.
+[2025], we use two types of probes: (1) Grid-wise inputs treat each square in the $10\times 10$ grid as a different data point, with inputs consisting of the $64$-dimensional LSTM state $(h,c)$ at a square, concatenated for the 3 layers.
+(2) Global inputs aggregate information from the entire 10x10 grid, resulting in a 6400-dimensional input for each layer at every timestep.
+The train and test dataset comprises states collected by evaluating the $\text{DRC}(3,3)$ on the hard Boxoban levels, excluding the first 5 steps of each episode as the plan is still forming.
+Bush et al.
+[2025] evaluated their probes on simple toy-levels similar to Appendix D, which don’t require long-term planning for most levels whereas we evaluate our probes on the hard set that can’t be solved greedily and thus require long-term plans.
+For evaluation of multi-class probes, F1 scores are computed as one-vs-all: the presence or absence of a particular class is the probe label.
+We search the best learning rate and L1 decay with grid-search by evaluating the F1 on a validation split of $20\%$ of timesteps from the hard levels (Table 1).
+Concepts (labels).
+We choose concepts that are likely to encode steps in the agent’s plan.
+Both the Agent-Directions and Box-Directions probes are grid-wise.
+These are central to our causality analysis and the same as Bush et al.
+[2025].
+We train five additional probes: Next-Box, Next-Target, Next-Action, Pacing, and Value, described in Appendix F.
+  • Agent-Directions probe.
+Predicts the direction the agent takes from a square $(x,y)$ at the nearest future timestep.
+It has 5 outputs: NV (No Visit), UP, DOWN, LEFT, RIGHT.
+The probe takes the 192-dim hidden state activations concatenated across the 3 layers at a square $(x,y)$ as input.
+If the agent visits the square in the future, the probe predicts the corresponding direction; otherwise, it predicts NV.
+Thus, for each image observation, we get a $10\times 10$ target.
+  • Boxes-Directions probe. Same, but predicts the direction any of the four boxes will move in at a given square.
+Causal intervention.
+To see if the NN uses these concepts in its algorithm, we edit the network’s activations to influence probe outputs [Li et al., 2023], and measure the causal effect: the fraction of cases where an intervention affects the agent’s action accordingly.
+[^4] This measurement is known in the literature as the average causal effect or average treatment effect [Holland, 1986], the expected difference in a variable (whether the direction changes) when a causal intervention is present or absent..
+For a linear probe on activations $h$ with parameter vector $p$, we intervene to increase the logit $p\cdot h$ to an adaptive value $\alpha$, following work on steering vectors [Turner et al., 2023, Rimsky et al., 2023, Li et al., 2024].
+More precisely, we find $\operatorname*{arg\,min}_{h^{\prime}}\|h^{\prime}-h\|_{2}^{2}$, with the constraint that $h^{\prime}\cdot p>\alpha$.
+The solution to this is $h^{\prime}=h+\hat{p}\max(0,\alpha-(h\cdot p))$, where $\hat{p}=p/\|p\|^{2}_{2}$.
+For multiclass direction probes, we also zero-out the logits of the other classes.
+For example, to influence the UP direction at a square, we set the UP logit to $\alpha$ and the DOWN, LEFT, RIGHT, and NV logits to 0.
+This intervention more precisely edits the network’s representation as compared to the intervention of $h^{\prime}=h+p$ performed by Bush et al.
+[2025] on toy-levels.
+Our intervention method is required to observe causal effects on standard difficult levels in the hard set.
+[…]
+Probe Pred (a) Causal Effect (b)
+F1 $\alpha$ Avg Best-case
+Box-Dir $86.4\pm 0.1$ 30 $49.3\pm 2.0$ $82.5\pm 2.5$
+Agent-Dir $72.3\pm 0.1$ 10 $7.1\pm 0.3$ $20.7\pm 0.7$
+Next box $74.2\pm 0.4$ 40 $5.5\pm 1.0$ $15.1\pm 2.5$
+Next target $54.3\pm 0.5$ 30 $4.6\pm 0.8$ $13.2\pm 2.0$
+3.2 Probe evaluation and causality
+Probe predictive power.
+Tables 1 and 10 show that most probes demonstrate high predictive accuracy.
+Figure 1 (right) and 17 show visualizations of Agent-Directions and Box-Directions probes [^5] The supplementary material provides visualization videos of all the probes across several levels..
+In particular, the box-directions probe can predict future box movements up to 50 steps in advance with over 90% accuracy (Figure 4, left).
+Only the Box-Directions probe is strongly causal.
+For each time step, we measure the causal impact of probes by intervening on the hidden state activations $h,c$ at the layer the probe was trained on.
+The results are reported in Table 1(b), where we perform a grid-search over the intervention strength $\alpha$ to find the optimal value.
+Despite our adaptive scaling, high values of $\alpha$ can still destabilize the agent’s behavior, causing random actions, while low values of $\alpha$ may fail to induce the intended behavior.
+Most probes show little causal effect, the Box-Directions probe (used in Figure 1) demonstrates strong causality, and the Agent-Directions probe shows mild causality.
+Even then, the Box-Directions probe seems to alter box movements only when they do not lead to naive deadlocks.
+For instance, the model avoids pushing a box into a wall if later steps indicate that it would need to get the box off the wall to get it to the target.
+To account for this behavior, we introduce the Best-case causal effect in Table 1(b).
+This tests all three alternate directions that a box could move, counting the probe as “causal” if it successfully influences any of these directions without causing a naive deadlock.
+By intervening on the probes, we can lock the DRC into a plan when there are two equally valued options.
+Figure 17 in Appendix D shows visualizations of how causal interventions with the directions probe alter the trajectory of the agent.
+However, in most cases, if we stop intervening with a suboptimal plan, the DRC formulates a better plan online and follows it.
+This and the naive deadlock problems may be caused by us intervening only on the current square of the box, target or agent: the NN activations still contain contradictory information about other squares, so the NN malfunctions.
+4 The plan improves with computation
+Guez et al.
+[2019] showed that the DRC solves more levels if we give it thinking steps at the start of episodes (see Definition 1.4).
+But what happens in the network during these thinking steps?
+Using the Box-Directions probe from Section 3, we show that the represented plan increases in length and accuracy (Figure 4, middle and right), demonstrating finding 2.
+Extra thinking is more helpful for levels that have longer optimal solutions (Figure 3, right), or which require less myopic thinking to place the first few boxes (Figure 3, left)(b).
+Thus, we are moderately confident that extra thinking improves the success rate by letting the network develop its plan long enough that it finds a non-myopic solution instead of accidentally locking the level.
 Figure 3: Left: Average time step to place each box $B_{i}$ on target for different numbers of thinking steps.
-(a) Averages across all levels where the box $B_{i}$ is placed on target by DRC with $N$ thinking steps.
-(b) Averages for all levels solved by $6$ thinking steps but not solved by 00 thinking steps.
-More thinking steps makes the DRC avoid greedy strategies in favor of long-term return.
+(a) Averages across all levels.
+(b) Averages over levels solved by $6$ thinking steps but not solved by 00 thinking steps.
+More thinking steps make the DRC avoid greedy strategies in favor of long-term return.
+The 95% confidence intervals computed with the bootstrap method over the levels is very small and not visible in the plot.
 Right: Average optimal solution length of levels grouped by the number of thinking steps at which the level is first solved.
@@
 NS stands for “not solved”.
-3.2 Agent “paces” to get more computation
+Effects of thinking time: non-myopic network
+We first examine how additional thinking time impacts the DRC performance by performing thinking steps at the start of an episode, allowing the network to process its hidden state.
+As shown by Guez et al.
+[2019], adding 6 thinking steps increases the success rate by $4.7\%$, with a slight drop beyond 16 steps (Figure 7).
+This improvement is more pronounced for harder levels, which require longer optimal solutions (Figure 3, right).
+Notably, thinking time does not correlate with the number of nodes expanded by an A* search, suggesting the DRC works differently (Figure 21).
+Thinking reduces early error in levels where placing the first box may block completion.
+Figure 3 (left) (b) shows that without extra thinking steps, the agent pushes the first box to a target about 4 steps too early, and notably much earlier than average (Figure 3, left).
+Providing 6 thinking steps prevents this mistake.
+While some improvement stems from avoiding catastrophic moves, the fact that the average time to push the first box exceeds 6 steps even without forced thinking suggests that the network uses the additional time to form a more complete solution, which changes behavior.
+The impact of thinking time varies with training duration and level difficulty.
+Figure 1 (bottom-left) shows that most of the planning improvement occurs within the first 70M steps.
+Beyond this point, the planning effect grows a bit more until 200M steps and then stabilizes (hard levels) or slowly decays (medium validation).
+Based on this, we speculate that the network develops better heuristics for when to think on medium-difficulty puzzles, reducing its default myopia.
+Plan improvement.
+If DRC refines its plan with extra computation (finding 2), probe-extracted plans should become more predictive of the agent’s actions as more computation time is provided.
+Figure 4 (right) confirms this effect: the plan initially has low accuracy and rapidly improves as DRC is given more thinking steps.
+Additionally, Figure 4 (middle) shows that the plans increase in complexity over time, with the length of the plans, defined as sum of length of continuous chains starting from boxes, growing significantly during these additional computation steps.
+Figure 4: Left: The number of steps in advance that the Box-Directions probe is able to predict a box move.
+The probe predicts $\geq 90\%$ of the actions 50 steps before they occur.
+Middle and Right: The plan quality, as measured by summing length of probe-predicted chain of box-directions starting from boxes and counting squares with positive non-empty predictions, and the F1-score increases over thinking steps (Section 3.1).
+This suggests that the DRC refines its plan during computation, including across the three ticks per environment step.
+5 Purposeful “pacing” to get more computation
 On occasion, the DRC exhibits a curious behavior: the agent “paces” in a cycle, returning to the same location, without touching any box.
-The per-step penalty makes this behavior naively sub-optimal, because it does not advance the puzzle state.
-Since the $\text{DRC}(3,3)$’s performance improves with thinking steps, could the agent be using pacing to gain computation time to improve its plan and so avoid irreversible actions which may block the solution?
-The following evidence suggests the agent has developed this meta-strategy:
-The agent paces mostly at the start of episodes.
-In some levels, a single sub-optimal step can lock the level, which makes calculating a valid plan in early steps beneficial.
-Figure 4 (left) shows most cycles start in the first 5 steps of the episode, and (middle) forcing the model to think for six steps eliminates about 75% of these early cycles.
-Thinking steps almost entirely substitute pacing
-We run the $\text{DRC}(3,3)$ on medium-validation levels and record when and where cycles occur. We merge all cycles which overlap with other cycles, for a total of 13 7021370213\,70213 702 cycles.
-We then make the DRC take $N$ thinking steps just before it would start a cycle of length $N$.
-As shown in Figure 4 (right), these thinking steps replace the cycles: comparing the cycle and the substitution, the DRC follows the same trajectory for at least $60\%$ of the levels and for a minimum of $30$ steps [^4] For context, the median solution length for train-unfiltered is exactly 30 steps..
-If our hypothesis that cycles are utilized for thinking holds true, we would expect to see uninterrupted periods of action after each cycle.
-In the $N$ steps after an $N$-step cycle concludes, $98.8\%$ of trajectories have no cycles.
-When substituting them for NOOPs, this becomes $82.4\%$, lower but still much higher than $0\%$; indicating that thinking steps largely remove the NN’s perceived need for cycles.
-Comparing $\text{DRC}(3,3)$ to ResNet
-It is possible that the DRC is pacing because it is not optimized enough.
-The ResNet is much worse than the $\text{DRC}(3,3)$ at Sokoban, but it is optimized enough to almost never perform cycles: since it has no memory, any cycles would make it stuck.
-So the amount of training is definitely enough to remove cycles if they have a big impact, and we believe it is likely to also remove cycles if they have a moderate negative impact.
-It could also be that the DRC is computing a multi-step heuristic (e.g.
-one that depends on the convolutional receptive field).
-Against this, the pool-and-inject layer in the DRC makes the receptive field less important for communication, and the ResNet (which performs worse) has a much larger receptive field per forward pass.
-Together, this is weak evidence that receptive-field heuristics are not the main mechanism.
-Heuristics may still play a role.
-For example, due to the -0.1 per-step penalty and discount rate, it is rational for the training process to make the RNN place boxes greedily (hence the increasing time-to-box in Figure 3, middle).
-It could be that the planning effect disappearing for medium levels (Figure 1, left) is due to better heuristics about when to be greedy and when to take time to think.
+This behavior does not advance the puzzle state, so it is sub-optimal due to the per-step penalty incurred.
+Why is it still present in the trained network?
+We hypothesize (3) that this behavior is incentivized by training, and its purpose is to give the DRC enough time to develop the plan.
+We show this with multiple lines of evidence, based on observing and intervening on NN behavior, and speed of change in the plan during cycles.
+We also re-train the $\text{DRC}(3,3)$ with smaller penalties for taking NOOP actions, and show that the total number of cycles stays constant while the number of NOOPs increases.
+Through all these experiments, we take care to highlight that most cycles are deliberate for plan computation, as opposed to a small fraction of cycles that could be occurring accidentally
+Unless otherwise noted, the following experiments are for cycles recorded on medium-validation levels. We merge overlapping cycles, resulting in a total of 13 7021370213\,70213 702 cycles.
+Cycles happen early in levels.
+In some levels, a single sub-optimal step can render the puzzle unsolvable, so it is very important to develop enough of the plan to prevent that.
+Section 4 showed that extra thinking steps help in large part by preventing these mishaps, so it makes sense that training would incentivize doing similarly early in a level.
+Accordingly, Figure 6 (left) shows that most cycles start in the first 5 steps of the episode.
+Plans improve much faster in cycles.
+If the DRC uses cycles to refine its plan, we expect the plans to be worse at the start of cycles (because they need improvement), and to change faster during the cycle.
+We check this by looking at the F1 score of box-directions probe at the first step of cycles, and the rate of change in F1 and plan length found with box-directions probe, as measured by number of positive predictions of the probe.
+Plan length and quality are highly influenced by whether the agent is close to the beginning or end of a level.
+To account for this, we pair each step in a cycle with a non-cycle step from a random episode, such that their timesteps $t$ match.
+Using the hard-level set, we find that the F1 score at cycle start points is $51.13\%\pm 1.52\%$, compared to $68.13\%\pm 1.45\%$ for non-cycles.
+Per step, the F1 score improves on average by $0.96\%\pm 0.15\%$ during cycles, versus $0.60\%\pm 0.15\%$ during non-cycles.
+Finally, the plan length (number of squares with a predicted action) grows by 2.03 squares per step during cycles, compared to 1.37 squares per step during non-cycles.
+Figure 23 shows the distribution corresponding to some of these aggregate statistics.
+While cycle-step improvements definitely skew larger, the distributions has some overlap, which suggests that some cycles could be accidental in which F1 improvement is not larger than non-cycles steps.
+Training with smaller NOOP penalty.
+The original version of the $\text{DRC}(3,3)$ could not take NOOPs, so it has to pace to get time to think.
+We train for 800M steps a version of the $\text{DRC}(3,3)$ which gets a smaller penalty for taking a NOOP than for moving.
+If the DRC is deliberately stalling when the plan is not good enough, we would expect it to use NOOPs when it is stalling deliberately.
+At the same time, the total number of cycles should stay constant: each step incurs a penalty, so it is still optimal to solve the level as quickly as possible.
+Figure 5 shows exactly this effect.
+The per-move penalty is fixed to $0.1$, and the NOOP penalty varies in $\{0.05,0.07,0.09,0.1\}$.
+We plot the total number per episode of steps in cycles (including NOOPs) and of NOOPs.
+As expected, the NOOPs per episode increase with smaller NOOP penalties, and the cycle-steps per episode stay roughly constant (slightly decrease).
+At the same time, the DRC(3,3) performance stays similar: it goes from solving $70\%$ to $60\%$ of levels.
+The trend of more cycles getting replaced with NOOPs suggest that most cycles are deliberate, but a small fraction could be accidental that occur anyways.
+We exclude cycles that happen after the last box is placed in a level because when the DRC sees a level is locked, it goes in cycles until the time limit, and that would make the results above largely depend on the time limit.
+Figure 5: Top: Average number of NOOPs and cycle steps per episode, when training with a NOOP action that has less penalty than the movement actions.
+This makes the network substitute many of the cyclic steps with NOOP steps on average, indicating it is a deliberate behavior.
+The total number of cycle steps (blue) stays constant.
+Bottom: percentage of validation-medium levels solved for each network.
+Error bars show the minimum and the maximum over 3 training seeds.
+Cycles can be substituted by NOOPs.
+Figure 6 (middle) shows that forcing the model to think for six steps eliminates about 75% of the early cycles.
+Thus, we can again conclude that most cycles are deliberately performed by the network.
+We can also directly replace would-be cycles with NOOPs.
+Just when the DRC would start a cycle of length $N$, we intervene and make it take $N$ thinking steps.
+Figure 6 (right) shows that these thinking steps largely replace the cycles: in at least $60\%$ of the levels, the DRC followed the same trajectory for a minimum of $30$ steps.
+For comparison, the median solution length for train-unfiltered is exactly 30 steps.
 Figure 6: Left: Histogram of cycle start times on the medium-difficulty validation levels.
-Middle: Total number of cycles the agent takes in the first 5 steps across all episodes in medium-difficulty validation levels with $N$ initial thinking steps.
-Right: on all levels, we replace $N$-length cycles with $N$ thinking steps and check the proportion of trajectories which are equal between these two treatments, for $x$ steps after the cycle.
-4 The $\text{DRC}(3,3)$ causally represents its plan
-In order to improve the plan with search, as the previous section suggests, the agent must have some internal representation of the plan.
-The absence of such representation would force us to discard the theory that the agent is doing search.
-This section presents probes which predict the future actions of the $\text{DRC}(3,3)$ and other various future features of the environment.
-Of these, the one which predicts the future direction boxes will move in has a strong causal effect in the actions of the RNN.
-Following the standard of evidence in Li et al.
-[2023], we declare this conclusive evidence that the $\text{DRC}(3,3)$ represents and uses plans.
-The spatial structure of the probes and some of the targets are due to Bush et al.
-[2024].
-We also observe the plan grow during thinking steps and, by intervening on the probes, are able to lock the RNN into a plan when there are two options of equal value.
-However in most cases, if we stop intervening on the RNN activations with a suboptimal plan, the RNN comes up with a better plan online and follows it.
-We have some evidence for the $\text{DRC}(3,3)$ considering multiple plans: for example, in the level in Figure 1 (right) the box probe first considers moving box 4 down and right (shown in red), but then rethinks that and ends up taking different actions.
-However, we have been unable to find multiple simultaneous representations of plans, or anything that would decide between them such as the value of the plans.
-In all, this section provides very little evidence in favor of search, but conclusive evidence for a causal plan.
-Figure 5: Success rate of the RNN, and plan quality as measured by box-direction probes (Section 4.1), all increase with thinking steps.
-We measure plan quality by 1) summing the length of chains of probe-predicted directions, which start in boxes, and 2) simply counting the number of squares in which the box-direction probe predicts something.
-The F1 score of the plan probe also increases over time, suggesting that the DRC builds the plan in that time.
-The right plot also includes plan quality evaluation on the three sequential ticks per environment step of the DRC(3, 3) network.
-4.1 Probe methodology: training, intervention and targets
-Training.
-We train logistic or linear regression probes with L1 decay to predict various features of the environment from the agent’s activations.
-We train on two types of inputs.
-1) Grid-wise inputs [Bush et al., 2024]: each square in the $10\times 10$ grid is a different data point, potentially with a different label depending on location and timestep.
-The input is the $64$-dimensional LSTM state $(h,c)$ at a square, usually concatenated for the 3 layers.
-2) Global inputs: Each time step produces only one data point.
-The probes take as input the 64 channels of all of the squares in the 10x10 grid, resulting in a 6400-dimensional input for each layer.
-In both cases, the per-layer activations are concatenated for probes that are trained on multiple layers.
-The dataset consists of states collected by evaluating the $\text{DRC}(3,3)$ on the hard Boxoban levels, excluding the first 5 steps of each episode because the plan is still forming.
-For multi-class probes, training and F1 scores are computed as one-vs-all: the presence of absence of a particular class is the probe label.
-We search the best learning rate and L1 decay with grid-search by evaluating the F1 on a validation split of $20\%$ of timesteps from the hard levels.
-We take the first 1000 medium-validation levels as our test set to report the results (Table 1).
-Causal intervention.
-Given a probe that is predictive of future behavior, we edit the agent’s activations such that the probe predicts the behavior we desire to induce.
-To the extent that the agent carries out the edited behavior, we can conclude that it executes a plan encoded in the linear direction of the probe.
-For a linear probe on the activations $h$ with parameter vector $p$, we intervene in order to increase the logits for a particular behavior, given by $p\cdot h$ (we omit a scalar intercept for brevity).
-Unlike previous work in steering vectors for LLMs [Turner et al., 2023, Rimsky et al., 2023, Li et al., 2024], we found that simply adding a perturbation equal to a fixed multiple of $p$ caused the policy to degenerate and choose random actions.
-We instead add an adaptively-scaled multiple of $p$ to the activations, chosen to minimize the change in the activations while still causing the probe’s logit to be greater than a desired amount $\alpha$.
-More precisely, we find $\operatorname*{arg\,min}_{h^{\prime}}\|h^{\prime}-h\|_{2}^{2}$, with the constraint that $h^{\prime}\cdot p>\alpha$.
-The solution to this is $h^{\prime}=h+{\hat{p}}\max(0,\alpha-(h\cdot p))$, where ${\hat{p}}=p/\|p\|^{2}_{2}$.
-Targets.
-We choose targets that plausibly encode the steps of a plan.
-Most of these labels are for grid-wise probes, but the pacing and value probes are global.
-The direction probes are very similar to Bush et al.
-[2024].
-  • Agent-Directions probe: This probe has 5 outputs: NV (No Visit), UP, DOWN, LEFT, RIGHT.
-The probe takes as input the 64-dim hidden state activations at a square $(x,y)$ at timestep $t_{i}$.
-If the agent visits the square $(x,y)$ at a future timestep $t>t_{i}$, then the probe predicts the corresponding direction that the agent takes from that square.
-For squares that are visited multiple times in the future, the direction corresponding to the first future visit is taken as the target.
-If a square is never visited in a future timestep, the probe predicts NV.
-There is one single multinomial probe, for all time steps and locations $(x,y)$.
-  • Boxes-Directions probe: This probe is similar to the above probe, except that it is trained to predict the direction in which any of the four boxes will move at a given square in the future.
-The rationale for this probe is that the boxes, and constraints on their movement given by obstacles, are the main difficulty with Sokoban puzzles.
-  • Next-Box probe: Predicts $1$ on the square of the box that the agent will move next and 00 for every other square.
-  • Next-Target probe: Predicts $1$ on the square of the target that the agent will put a box in next, and 00 for every other square.
-  • Pacing probe: The global label is 1 if the agent is currently in a cycle, and 0 otherwise.
-  • Value probe: The global label is the numerical value that the critic head outputs.
-[…]
-Probe target Predictive Probe Results (a) Causal Probe Results (b)
-Sparse Probes Causal Effect
-Best F1 F1 $L_{0}$-norm $\alpha$ Average Best-case
-Box-directions $86.4\pm 0.1$ $73.1\pm 0.1$ $63$ 30 $43.7\pm 0.6$ $77.8\pm 0.9$
-Agent-directions $72.3\pm 0.1$ $61.3\pm 0.1$ $121$ 10 $7.1\pm 0.3$ $20.7\pm 0.7$
-Next box $74.2\pm 0.4$ $69.7\pm 0.5$ $51$ 40 $5.5\pm 1.0$ $15.1\pm 2.5$
-Next target $54.3\pm 0.5$ $44.0\pm 0.5$ $32$ 30 $4.6\pm 0.8$ $13.2\pm 2.0$
-Pacing $31.0\pm 1.8$ $31.0\pm 1.5$ $5$
-4.2 Probe evaluation and causality
-Probe predictive power.
-Most probes are quite predictive, as seen in Table 1.
-The exception is the pacing probe, with $F_{1}=31.0\%$, which is not much better than the constant 1 probe, which has $F_{1}=12.8\%$.
-This leads us to conclude that the $\text{DRC}(3,3)$ does not represent whether it is in a cycle or not.
-See Figure 1 (right) and Figure 6 for visualizations of agent-direction and box-direction probes [^5] The supplementary material provides visualization videos of all the probes across several levels..
-For the value function probe, we compute the fraction of variance explained $R^{2}$.
-If we train a global probe, $R^{2}$ is very high: $97.7\%$$99.7\%$ depending on the layer.
-However, grid-wise probes obtain much worse but still passable results: $41.0\%$$79.2\%$ depending on layer.
-We visually checked whether the grid-wise probe reads off the values of different plans, but could not find any such pattern.
-Almost all the performance of the global probe is recovered by training on the mean-pooled inputs: $95.2\%$$99.5\%$.
-It is likely that the global and mean-pooled value probes are indirectly counting the number of squares the agent will step on, which almost fully determines the value, and we know is possible due to the future-direction probes.
-Only the Box-Directions probe is strongly causal.
-We measure how causal the probes are by intervening at every step on the hidden state $h,c$ at the layer the probe was trained on, following the procedure in section 4.1 and checking whether the agent follows that action.
-We report the result in Table 1(b) by performing a grid-search over the strength of intervention $\alpha$ and picking the best value.
-Despite our adaptive scaling of the intervention, a high value of $\alpha$ can still disturb the agent such that it starts performing random actions whereas a low value of $\alpha$ may not be sufficient to cause the agent to follow the intended behavior.
-Most of the probes are not causal: only the “boxes directions” probe (used in Figure 1) is causal while the “agent-direction” probe mildly causal.
-Even then, we have observed that it can only change boxes directions when they would not deadlock the level in a naive way.
-That is, if you try to make the model push a box onto a wall, such that the box would not be able to reach a target anymore, it does not.
-For this reason, we introduce the Best-case causal effect in Table 1(b): try all three directions that are not the actual direction that the box follows, and count the probe as “causal” if it works for any of them (e.g.
-because they do not naively deadlock the level).
-The box-direction probe’s coverage increases during thinking time.
-Figure 5 (middle) show that the amount of squares in which the probe predicts a direction, be they lone or linked to a box, grows over thinking steps.
-Its description of what the box movements will be also greatly improves after a few steps of thinking Figure 5 (right).
-4.3 Bistable and unstable plans in toy environments
-The causal intervention results from Table 1 show that the box-directions probe is much more causal than the agent-directions probe.
-We further study this in an isolated setting.
-Figure 6 (left) shows a level with only a single box next to a target, where the agent has only two equally good paths it can follow.
-Without intervening, the agent first takes the path on the right by moving three steps, while the agent-direction probe grows towards the box.
-After the first three moves, the chain connects.
-The agent then takes a U-turn to follow the path on left.
-This is very unlike the optimal behavior, which is to commit to any path and follow it.
-We can get the agent to commit to one path early by intervening with the agent-direction probe on any of the two paths on the first step only, as shown in Figure 6 (middle).
-When we construct a similar level where there are two equally good paths for a box to follow, we find that the agent-direction probe is no longer causal; but the boxes direction probe does get the agent to commit.
-Figure 6 (right) shows an almost empty level.
-When we intervene with the box-directions probe on the first four steps, the agent follows the laid out path in those steps, but then quickly switches to a more optimal path with fewer turns.
-The laid out path (not taken, shown in red) is not followed after stopping the intervention.
-This analysis provides strong evidence that the network primarily builds upon the boxes directions represented in the activations and relies on the agent-direction when the boxes directions do not inform the subsequent actions.
-Figure 6: Left: A custom level where the agent has two equally good paths to follow.
-The arrows show the prediction of the agent-direction probe with opacity proportional to the number of times an arrow was predicted across all the steps.
-The green and red arrows are correct and incorrect predictions, respectively.
-The agent behaves suboptimally by returning to the start and going left after first going right for three steps.
-Middle: The agent takes the path on the right after intervening with the corresponding arrows using the agent-direction probe on the first step.
-The same happens on the left if we intervene on that path without the suboptimal steps right of the undisturbed agent.
-Right: An empty level with box-directions probe intervening on the first four steps which are followed correctly by the agent on those four steps.
-When the intervention is removed on the fifth step, the agent computes the simpler path in green and doesn’t follow the path laid out earlier (in red).
-5 Model surgery exposes latent generalization capabilities
-A common feature of search algorithms is that they generalize very well to unusual problems in their domain.
-This was the main thrust of the argument of Guez et al.
-[2019]: they exhibited many ways in which the $\text{DRC}(3,3)$ generalized to levels outside of the training distribution to argue that it does search.
-5.1 Generalizing beyond $10\times 10$ inputs and training examples
-This section goes beyond the capabilities of $\text{DRC}(3,3)$ as a black-box unit.
-By looking for interpretable features in the activations of the ConvLSTM core, we found that layer 3, at the last tick, has four channels which represent the next action to take.
-Since the network up until the ConvLSTM core is completely convolutional, we can evaluate it on inputs of any size – unlike $\text{DRC}(3,3)$ as a whole, which only works on $10\times 10$ inputs, and is trained only on levels with four boxes.
+Middle: Total cycles the agent takes in the first 5 steps across all episodes with $N$ initial thinking steps.
+Right: Proportion of trajectories that match when replacing $N$-length cycles with $N$ thinking steps, tracked for $x$ steps post-cycle.
+6 Generalizing beyond $10\times 10$ inputs and training examples
+We extend the DRC’s capabilities by modifying its internal structure to generalize beyond the original $10\times 10$ input size and four-box training distribution.
+Analysis of the ConvLSTM core reveals that layer 3, at the final tick, encodes four key channels representing the next action provided in Table 10.
+Since the ConvLSTM is completely convolutional, it can process inputs of arbitrary sizes, unlike the full DRC architecture, which flattens the convolution output and passes through an MLP layer.
 Spatial aggregation.
-We start with the Layer 3 grid-wise next-action probes from Table 3.
-For each spatial location, these predict the action the $\text{DRC}(3,3)$ will take in this step.
-We attempt three different spatial aggregation methods: mean-pooling, max-pooling, and the proportion of locations with positive probe readouts for a particular action.
-We learn their relative weights and a baseline for each action by using Adam with learning rate $\ell=10^{-3}$, annealed linearly to 0 for 10000 steps.
-Table 2 lists the learned parameters, which place most weight on the mean-pooling aggregation rule.
-The training set consists of 3000 levels, 1000 from each of the Boxoban unfiltered, medium, and hard training sets.
-Table 2: Weight and bias for transforming direction probes to predictions
-Mean Max Positive proportion
-Weight 1.2086 -0.0582 0.2070
-Up Down Left Right
-Bias 0.3337 -0.0921 -0.0632 -0.0539
+Using layer 3 next-action probes (Table 10), we predict at each spatial location the next action of the DRC.
+To turn the probe’s grid-wise outputs into a single action prediction, we need to aggregate them spatially.
+We linearly combine three spatial aggregation methods: mean-pooling, max-pooling, and the proportion of positive probe outputs.
+The optimal combination weights and a bias for each action are learned by minimizing cross-entropy loss against the MLP block using the Adam optimizer on 3000 levels (1000 from each Boxoban training set).
 Results.
-The aggregated features obtain 83% accuracy at predicting the MLP’s action output on the training set, which is less than expected given the $90\%+$ F1 scores of probes in Table 3.
-With only seven parameters, we did not bother collecting an i.i.d.
-test set and went straight to the generalization tests.
-The adjusted ConvLSTM is able to solve many levels that are out-of-distribution: larger than $10\times 10$ in both dimensions, and with more than four boxes, e.g.
-those in Figure 2.
-Out of our test set collected from Þorsteinsson [2009], it can solve $59/480$ levels with both dimensions $>10$, $30/203$ levels with only one dimension $>10$ and $180/482$ levels with both dimensions $\leq 10$.
-Appendix B has a breakdown by level collection.
-5.2 Looking for interpretable features with Sparse Autoencoders
-The way we found out actions were represented in the ConvLSTM was with sparse autoencoders (SAEs).
-We inspected an SAE [Huben et al., 2023, Bricken et al., 2023] trained on the last layer with the top-k activation [Gao et al., 2024, Tamkin et al., 2024], $k=8$.
-However, we then noticed that the action features were also represented directly in the ConvLSTM channels.
-Table 3 compares raw channels, SAE features, and linear probes for how well they match the interpretation of “the feature fires when the $\text{DRC}(3,3)$ takes a particular action”.
-We find that the raw channels and linear probes are similarly good, and both are considerably better than the SAE.
-Thus, for these features, SAEs actually worsened interpretability. We also found three other interpretable features in the SAE, which were also represented directly in the channels (Appendix E).
-Table 3: Action features scores across channels, probes, and SAE features
-Feature Channel SAE Feature Probe
-Number Prec Rec F1 Number Prec Rec F1 Prec Rec F1
-Up L3C29 95.7 88.1 91.7 L3F270 93.9 76.2 84.1 97.5 86.5 91.7
-Down L3C8 98.4 80.8 88.8 L3F187 98.0 79.1 87.6 97.6 86.9 91.9
-Left L3C27 85.5 84.6 85.1 L3F244 96.1 63.2 76.2 83.5 86.6 85.0
-Right L3C3 97.0 86.9 91.7 L3F385 94.6 78.5 85.8 97.6 87.4 92.2
+The aggregated features achieve $83\%$ accuracy on the training set and $77.9\%$ on medium-difficulty validation levels.
+Although lower than individual probes F1 scores Table 10, these enable the adjusted ConvLSTM to solve many out-of-distribution levels, including larger grids and more than four boxes, e.g., Figure 2 (right) and Figure 15.
+Specifically, from the test set by Þorsteinsson [2009], it solves $64/484$ levels with both dimensions $>10$, $38/203$ levels with only one dimension $>10$, and $221/486$ levels within the original constraints.
+A detailed breakdown by level collection of performance and largest level solved is provided in Appendix C.
+The largest level solved per level-collection is between 3-4x larger in area and number of boxes compared to the $10\times 10$ levels with 4 boxes seen during training.
+Our findings show that by aggregating interpretable spatial features, the ConvLSTM generalizes to larger and more complex puzzles without retraining.
+This suggests that much of the DRC’s planning ability is encoded in the ConvLSTM core, and refining spatial predictions allows for scalable generalization in neural planning systems.
+Stress-testing on zig-zag.
+How well does this extension to larger levels work?
+We stress the network by placing it in straightforward (but long) $n\times n$ zig-zag levels (Figure 16).
+The $\text{DRC}(3,3)$ solves all levels for $n\leq 15$, but fails at $n\geq 16$ no matter how many thinking steps.
 7 Related Work
-Interpreting agents and planning.
-Several works have attempted to find the mechanism by which a simple neural network does planning in mazes [Mini et al., 2023, Knutson et al., 2024, Brinkmann et al., 2024], gridworlds [Bloom and Colognese, 2023], and graph search [Ivanitskiy et al., 2023].
-Men et al.
-[2024] investigates LLM reasoning in a simple block-stacking task.
-We believe the DRC we present is a clearer example of an agent than what these works focus on, and should be similarly possible to interpret.
-Concurrent work [Bush et al., 2024] interprets another DRC Sokoban agent and can also predict future actions. That team invented grid-wise probes and the Agent-Directions and Box-Directions targets.
-Other works interpret superhuman game-playing agents.
-Jenner et al.
-[2024] find evidence of lookahead in Leela Chess Zero [Team, 2018], without predicting future actions.
-McGrath et al.
-[2021], Schut et al.
-[2023] find interpretable concepts in AlphaZero [Silver et al., 2017].
-Probing for world models.
-Planning usually requires a model of the world.
-Li et al.
-[2023], Nanda et al.
-[2023b], Karvonen [2024] examine agents trained to predict possible moves in games, and are able to probe for the board state.
-Wijmans et al.
-[2023] also probe the state of navigation neural networks.
-Gurnee and Tegmark [2023] find a representation of positions in Earth for location-associated tokens in an LLM.
-Goal misgeneralization and mesa-optimizers for alignment.
-From the alignment perspective, AIs optimizing monomaniacally for a goal have long been a concern (e.g.
-Yudkowsky, 2006, Omohundro, 2008; see the preface of Russell, 2019).
-In a machine learning paradigm [Hubinger et al., 2019], the goal of the training system is not necessarily optimized; instead, the NN may optimize for a related or different goal [Di Langosco et al., 2022, Shah et al., 2022] or for no goal at all.
-Chain-of-thought faithfulness.
-Large language models use chain of thought, but are they faithful to it, or do they think about their future actions in other ways [Lanham et al., 2023, Pfau et al., 2024]?
-One could hope that LLMs perform all long-term reasoning in plain English, allowing unintended human consequences to be easily monitored, as in Scheurer et al.
-[2023].
-Ethical treatment of AIs.
-Do AIs deserve moral consideration?
-Schwitzgebel and Garza [2015] argue that very human-like AIs are conceivable and clearly deserve rights.
-Tomasik [2015] suggests that most AIs deserve at least a little consideration, like biological organisms of any species [Singer, 2004].
-But what does it mean to treat an AI ethically?
-Daswani and Leike [2015] argue that the way to measure pleasure and pain in a reinforcement learner is not by its absolute amount of return, but rather by the temporal difference (TD) error: the difference between its expectations and the actual return it obtained.
-If the internals of the NN have a potentially different objective [Hubinger et al., 2019, Di Langosco et al., 2022], then the TD error should come from a place other than the critic head.
-Research into learned search algorithms is an early step toward finding the learned-reward internal TD error, if it exists.
-This could be a higher-assurance complement to simply asking the AI [Perez and Long, 2023].
-Fully reverse engineering small networks.
-Many works reverse engineer all of, or most of, a small NN that does an algorithmic task [Nanda et al., 2023a, Chughtai et al., 2023, Zhong et al., 2023, Quirke and Barez, 2023].
-Neural network architectures that reason.
-Many papers try to enhance NN thinking by altering the training setup or architecture.
-Schwarzschild et al.
-[2021b, a], Bansal et al.
-[2022] improve and evaluate a specific kind of RNN training that produces algorithmic thinking, though Knutson et al.
-[2024] argue that they do not generalize enough and thus do not implement the correct algorithm.
-Other works endow RNNs with variable amounts of computation per step, with [Chung et al., 2024] or without [Graves, 2016] an explicit world model.
+Our work contributes to the growing field of mechanistic interpretability focused on understanding reasoning and planning processes within neural networks.
+While the internal mechanisms of complex agents remain largely opaque, several lines of research provide context for our investigation.
+Interpreting planning and reasoning in neural networks.
+Prior studies have investigated planning mechanisms in simpler settings like mazes [Mini et al., 2023, Knutson et al., 2024, Brinkmann et al., 2024], gridworlds [Bloom and Colognese, 2023], and graph search problems [Ivanitskiy et al., 2023].
+Others have explored reasoning in Large Language Models (LLMs) on tasks such as block-stacking [Men et al., 2024].
+These works often focus on identifying representations of state or specific concepts, rather than studying the representation involved in planning.
+Researchers have also investigated sophisticated game-playing agents like Leela Chess Zero [Jenner et al., 2024] and AlphaZero [McGrath et al., 2021, Schut et al., 2023], uncovering evidence for lookahead mechanisms or high-level local concepts important for the game.
+Our work builds on these efforts but specifically targets the explicit representation and causal role of plans (complete sequences of future actions, Definition 1.1) within the DRC network trained on Sokoban, a game demanding complex, foresightful planning.
+Sokoban planning and the DRC agent.
+The DRC network trained on Sokoban provides a compelling case study of long-term deliberate planning owing to its ability to utilize extra thinking time [Guez et al., 2019].
+Our work aims to understand and explain this behavior by interpreting the DRC’s internal states.
+Concurrently with our research, Bush et al.
+[2025] also interpreted a DRC agent trained on Sokoban, by introducing probes similar to ours that predict future box and agent movements and found evidence for plan representations (similar to our finding 1).
+Our work complements and extends theirs in several key ways.
+Bush et al.
+[2025] performed causal interventions on $200$ handcrafted toy levels, whereas we perform more rigorous causal tests on $24k$ datapoints sampled from difficult standard benchmark levels and devise an intervention methodology that edits the representations more precisely (Section 3).
+Additionally, our work yields novel findings: finding 4 offers further evidence from probe-guided OOD generalization that DRC’s planning capability generalizes to larger levels not supported by the architecture, while finding 3 on pacing behavior provides explanation for the emergence of the performance improvements through “thinking time”.
+Probing and intervention methods.
+Our methodology relies on probing internal activations to decode information embedded linearly and intervening on these activations to test causality.
+Linear probing has been widely successful in finding representations of game states [Li et al., 2023, Nanda et al., 2023b, Karvonen, 2024] or spatial information [Wijmans et al., 2023] in various networks trained on game sequences.
+Adaptive computation for reasoning.
+The finding that the DRC’s plan improves with computation (Finding 2) and that the agent learns to "pace" (Finding 3) connects to research on improving neural network reasoning by altering training setup or architecture and leveraging adaptive computation time in RNNs [Schwarzschild et al., 2021b, a, Bansal et al., 2022].
+[Graves, 2016] and models that explicitly vary computation [Chung et al., 2024].
+However, Knutson et al.
+[2024] argue that they lack generalization and fail to implement correct algorithms.
+Other works explore models that adjust computation per step, either with an explicit world model [Chung et al., 2024] or without one [Graves, 2016].
+Our work suggests such adaptive strategies for long-term planning and reasoning can be learned implicitly using model-free RL, mirroring the recent advancements of LLMs leveraging test-time compute for improved reasoning [DeepSeek-AI et al., 2025, OpenAI, 2024].
+Goal misgeneralization and mesa-optimization.
+Understanding how neural networks perform tasks requiring sequential reasoning is crucial for ensuring transparency and safety of AI systems.
+In AI alignment, goal misgeneralization occurs when a neural network optimizes for unintended proxy objectives instead of its intended goal (e.g.
+Yudkowsky, 2006, Omohundro, 2008; see Russell, 2019).
+This risk relates to mesa-optimization, where a network develops internal objectives that diverge from training targets [Hubinger et al., 2019].
+Recent work has shown that networks may pursue goals misaligned—or even directly opposed—to those intended by their designers [Di Langosco et al., 2022, Shah et al., 2022].
+Mechanistic interpretability of AI agents is a crucial tool for understanding how neural networks internally represent goals and plans and for identifying and fixing potential misalignments.
+We discuss more related work in Appendix I.
 8 Conclusion
-We replicated a small RL-based Sokoban agent, whose performance benefits from increased compute in the form of externally imposed no-op steps [Guez et al., 2019].
-We find that the agent learns to explicitly exploit this capability by pacing, which substitutes for the external no-op thinking steps.
-Extra computation helps with longer levels and those that require non-myopic thinking.
-We show that the agent causally represents plans (sequences of actions to take).
-The most fundamental plan representation we find is the box-direction probes, which can predict and control where the agent pushes a box.
-By building up a chain of box probe steps over time, the agent is able to figure out how to place boxes to solve the level.
-Secondarily, we find agent-direction probes that take lower precedence than the box-direction probes, but can control the agent moves in isolated setups.
-We additionally find that the box probe chains get longer and longer with more thinking steps (Figure 5).
-Finally, we find that extracting a decision from the plan takes place almost entirely within the ConvLSTM, with the MLP doing little more than decoding it.
-To demonstrate this, we show that the agent can generalize to levels bigger than what it had ever seen during training by manually replacing its MLP decoder based on our mechanistic understanding.
-We hope the model organism we trained and open-source will catalyze research into NNs that learn to plan, and to understand their inner objectives.
-Acknowledgements
-We would like to thank ChengCheng Tan and Siao Si Loo for editing this paper, and the rest of the FAR team for general support.
-We also thank Alex F.
-Spies and anonymous reviewers for improving the paper with their comments and edits.
-We are also very grateful for the many questions about their RL setup that Arthur Guez and Timothy Lillicrap answered, as well as inspiration from Stephen Chung and Scott Goodfriend.
-Reproducibility Statement
-We provide the code and the trained models in the supplementary materials.
-We open-source our code and trained model(s).
-Additionally, sections 2, 4.1 and appendices A, D, and E describe the specific experimental details and settings we used to carry out our experiments.
+Building on prior and concurrent work on search-like behaviors of DRC agents [Guez et al., 2019] and plan representations [Bush et al., 2025], we provide detailed evidence that DRC agents represent causal plans through spatial encodings of future box movements and refine them over time.
+We describe how extra thinking steps increase success rate, by preventing greedy actions before the plan stabilizes into a longer-term solution.
+We find the agent takes advantage of this by “pacing” on-distribution to get enough time to stabilize the plan.
+By substituting the network output module with much simpler probes, we demonstrate how a mechanistic understanding can enable generalization beyond the agent’s training distribution.
+These findings advance the understanding of planning dynamics in neural networks and lay the groundwork for more transparent, safer and interpretable AI systems.
+Impact Statement
+This research into interpretability can make models more transparent, which helps in making models predictable, easier to debug and ensure they conform to specifications.
+Specifically, we train and open-source a model organism which is planning, and analyze it somewhat; we hope this will catalyze further research on identifying, evaluating and understanding what goal a model has.
+We hope that directly identifying a model’s goal lets us monitor for and correct goal misgeneralization [Di Langosco et al., 2022].
+Goal evaluation also has implications for AI welfare.
+Knowing the goals of a model would also make it possible to know whether these goals are being satisfied enough, and thus perhaps give a way to evaluate whether or not a model is happy (Appendix I), assuming it is a moral patient.
 Appendix A Training the test subject
@@
 A.1 Training hyperparameters
-All networks were trained with the same hyperparameters, which were tuned on a combination of the ResNet and the $\text{DRC}(3,3)$.
-These are almost exactly the same as Guez et al.
-[2019], allowing for taking the mean of the per-step loss instead of the sum.
+All networks were trained with the same hyperparameters, tuned on both ResNet and the $\text{DRC}(3,3)$.
+These largely match Guez et al.
+[2019], except we take the mean per-step losses instead of summing.
 Time limits.
@@
 A.3 Learning curve comparison
-It is difficult to fully replicate the results by Guez et al.
-[2019].
+Replicating the results of Guez et al.
+[2019] proved challenging.
 Chung et al.
@@
 [2019] in Figure 8, which shows that the learning curves for $\text{DRC}(3,3)$ and ResNet are compatible, but not the one for DRC(1,1).
-Our implementation also appears much less stable, with large error bars and large oscillations over time.
-We leave addressing that to future work.
-The success rate in Figure 7 is computed over 1024 random levels, unlike the main body of the paper.
+Our implementation exhibits reduced stablility, with large error bars and pronounced oscillations over time.
+We defer addressing this to future work.
 Table 3 reports test and validation performance for the DRC and ResNet seeds which we picked for the paper body.
@@
 ResNet: 3,068,421 (3.07M)
+Figure 7: Success rate of $\text{DRC}(3,3)$ on Validation-medium levels across forced thinking steps at the start of the episode. Increasing thinking steps increases performance.
 Figure 8: Success rate for Test-unfiltered and Validation-medium levels vs.
@@
 The dotted lines are data for the performance of architectures extracted from the [Guez et al., 2019] PDF file.
-The values are slightly different from what Figures 5 and 3 report because they are calculated on a random sample of 1024 levels (24 levels are repeated for test-unfiltered, which only has 1000 levels).
 Table 3: Success rate and return of DRC and ResNet on the unfiltered test set at various training environment steps.
@@
 2B 97.9 10.29 99.3 10.52 59.4 1.16 76.6 4.52
-Appendix B Generalizing the $\text{DRC}(3,3)$ to larger levels
+Appendix B Results on $\text{DRC}(1,1)$ and ResNet
+B.1 $\text{DRC}(1,1)$
+In this section, we show that the $\text{DRC}(1,1)$ network, which only has a single recurrent convolutional block that runs a single forward pass per environment step, also exhibits similar planning behavior to the $\text{DRC}(3,3)$ network, although achieving lower performance as expected due to lower parameter count and compute expenditure per environment step.
+The parameter count for the $\text{DRC}(1,1)$ and other networks used in the paper are reported in Table 2.
+Figure 9 shows the performance of the $\text{DRC}(1,1)$ network on the medium difficulty validation levels during training.
+The $\text{DRC}(1,1)$ network shows a similar planning effect as the $\text{DRC}(3,3)$ network (Figure 1), which matches the results of Guez et al.
+[2019] who also showed that extra thinking steps improved performance of both the $\text{DRC}(1,1)$ and the $\text{DRC}(3,3)$ network.
+We train the box-directions probe on the $\text{DRC}(1,1)$ network and evaluate it on the medium validation levels.
+The resulting probe achieves an F1 score of $72.3\%$.
+The F1 score is high enough to suggest that the $\text{DRC}(1,1)$ network also has a plan representation similar to the $\text{DRC}(3,3)$ network.
+Upon inspecting the probe visualizations, we find that the probe is indeed good at predicting the box movements in the short-term but is worse than the probe on $\text{DRC}(3,3)$ at predicting the long-term box movements (Figure 12, left).
+This suggests that the $\text{DRC}(1,1)$ network has a plan-representation, but it is limited to short-term planning.
+This results matches the evidence from the network’s overall performance which is significantly worse on level requiring long-term planning.
+We now ask whether the performance improvement during thinking steps can be explained by the network refining its plan as we showed for the $\text{DRC}(3,3)$ network in Section 4.
+We measure the plan quality as we did in the main text using the length of plan from boxes and number of positive predictions.
+We find that the plan quality (Figure 12, middle) as well as the F1 score (Figure 12, right) measured using the probe increases over the number of thinking steps.
+Figure 11 (left) shows that this network also avoids myopic strategies when given extra thinking steps.
+This shows that the $\text{DRC}(1,1)$ network also refines its plan during the thinking steps.
+Finally, we show that the $\text{DRC}(1,1)$ network also exhibits the pacing behavior as the $\text{DRC}(3,3)$ network.
+Figure 10 (left) shows that $\text{DRC}(1,1)$ also performs cycles at the start of levels, where extra computation steps are more useful to come up with better solutions.
+Figure 10 (right) shows that 60% of these cycles disappear with extra thinking steps, suggesting that they are performed by deliberately and not accidentally.
+The percentage is lower than 75% for $\text{DRC}(3,3)$ suggesting that $\text{DRC}(1,1)$ performs comparatively more accidental cycles, which is again consistent with lower planning-capacity and performance of the network.
+Hence, the main Findings 1--3 analysed for $\text{DRC}(3,3)$ in the main text are also true for $\text{DRC}(1,1)$, with results suggesting worse long-term planning as compared to $\text{DRC}(3,3)$ as expected.
+Figure 9: The $\text{DRC}(1,1)$ network shows similar planning effect as the $\text{DRC}(3,3)$.
+The plot is computed on the medium difficulty validation levels.
+The planning effect is measured as the difference in performance between the best thinking step and no thinking step.
+It achieves a lower performance than the $9$ layer ResNet network due to lower parameter count.
+Figure 10: Left: $\text{DRC}(1,1)$ also performs cycles at the start of levels when it is more beneficial.
+Right: $\text{DRC}(1,1)$ performs many cycles when given no thinking steps.
+More than 60% of the cycles disappear with extra thinking steps.
+Figure 11: Left: Thinking steps also help $\text{DRC}(1,1)$ to avoid myopic plans, thus increasing the time to place boxes first 3 boxes in solved levels while solving the level faster by placing the box 4 earlier.
+Right: The average optimal length of level solved by extra thinking steps doesn’t change with more thinking steps.
+This suggests that, unlike $\text{DRC}(3,3)$ which solves more difficult levels with more thinking steps, $\text{DRC}(1,1)$’s planning capacity is limited.
+Figure 12: Left: For $\text{DRC}(1,1)$, the number of steps in advance that the Box-Directions probe is able to predict a box move shows similar trend as $\text{DRC}(3,3)$, but is overall lower, indicating that the $\text{DRC}(1,1)$ is less capable of planning.
+Middle and Right: The plan quality, as measured by summing probe-predicted chain lengths starting from boxes and counting squares with positive non-empty predictions, and the F1-score increases over thinking steps (Section 3.1).
+This suggests that $\text{DRC}(1,1)$ also refines its plan during computation, although it has an overall lesser F1 score of $60\%$ as compared to $84\%$ for $\text{DRC}(3,3)$.
+Figure 13: Left: The box-directions probe trained on all layers for the ResNet model can predict moves with $80\%$ accuracy 50 steps before they occur.
+This implies that the ResNet model also computes a long-term plan for every observation.
+Middle: The plan quality and accuracy measured with the all-layer box-directions probe doesn’t improve with more environment steps as expected, since the plan in the activations need to be recomputed entirely from scratch on every step.
+This is unlike the DRC models which improve their plans with more thinking and environment steps.
+Right: Training the box-directions probe on each layer separately reveals that the plan representation is computed and improved in the early middle layers 1-4 with F1-score reaching 63.9%.
+The layer likely compute different parts of the plan since the probe trained on all the layers achieves a significantly higher F1-score of 84.5%.
+B.2 ResNet
+We now check if our findings about the DRC networks also generalize to a ResNet model with no recurrence over hidden state.
+The training and architecture details of the ResNet model is provided in Appendix A.
+We first train the box-directions probe on the concatenation of the output of the ReLU activations from every residual block in every layer.
+This creates a single input vector of size $1024$ for the linear probe.
+The activations are collected on 1000 uniformly sampled levels from the medium-difficulty train set on which the ResNet model has a solve rate of $59.9\%$.
+The resulting probe achieves an F1 score of $84.5\%$ for predicting future box-movement directions on the validation set, similar to the probe on the $\text{DRC}(3,3)$ network.
+Figure 13 (left) also shows that the activations of the model in early steps are highly predictive of the actions it will take many steps in the future.
+This suggests that both the networks have similar representations that store their long-term plans.
+However, since the ResNet model achieves a lower performance on medium and hard difficulty sets, we can conclude that the ResNet model is not as good at constructing correct plans as the $\text{DRC}(3,3)$ network.
+We found that plan improves with more thinking steps for the DRC model.
+Since the ResNet is not a recurrent model, it cannot take in extra thinking steps like the DRC networks can.
+Since the ResNet does not have a recurrent hidden state, we would expect that the plan representation of the model should not improve with more environments steps, as the activations are recomputed from scratch on every step.
+Figure 13 (middle plots) show that the plan is computed in one-step, and more environment steps do not result in better plan representations.
+Figure 13 (right) shows the validation F1-score of the same probe trained on the activations of the individual layers separately.
+We can see that the plan improves in the early to middle layers, implying that the plan representations are sequentially through the layers.
+Since the ResNet is state-less, it cannot exhibit the pacing-behavior.
+Hence, we can conclude that only the finding 1 of plan representation holds for the ResNet model, with plan-improvements happening sequentially through the layers.
+Appendix C Generalizing the $\text{DRC}(3,3)$ to larger levels
 We license the levels by Þorsteinsson [2009] as GPLv3, and make them available in the supplementary material.
-The performance of the $\text{DRC}(3,3)$ on each set of levels is in Table 6.
-The sets with higher scores are also those which humans find easier: for example, “Dimitri & Yorick” was made for children by Jacques Duthen and consists entirely of relatively small levels (largest is $12\times 10$) with at most 5 boxes.
-The sets where the $\text{DRC}(3,3)$ solves nothing are also difficult for the authors of this paper.
-We tried giving the $\text{DRC}(3,3)$ from 2 to 128 extra thinking steps for these levels, incrementing in powers of two.
+Table 4 shows the performance of DRC acriss various level sets.
+Higher scores align with levels considered easier by humans; for example, “Dimitri & Yorick” was made for children by Jacques Duthen, features small levels (maximum size: $12\times 10$) with at most 5 boxes.
+Level sets where the DRC fails to solve any puzzles were also challenging for the authors.
+We tested the $\text{DRC}(3,3)$ with 2 to 128 extra thinking steps, incrementing in powers of two.
 For most sets, we find some benefit to 2-4 thinking steps, but no more.
-The sole exception is XSokoban, which contains one level (Figure 2, right) which requires 128 steps of thinking to solve.
+The sole exception is XSokoban, which contains one level (Figure 15(c)) requiring 128 steps of thinking to solve.
 We encourage the reader to go to the website by Þorsteinsson [2009] and try solving the levels.
@@
 Table 4: Performance of the $\text{DRC}(3,3)$ on each set of levels by [Þorsteinsson, 2009].
-The “Max solved” columns represent the proportion of levels solved at the number of steps in the “at steps” column, which is the highest solved proportion for each number of thinking steps tried.
-All levels Levels larger than $10\times 10$
-Level collection # solved Max slv. max at # solved Max slv. max at
-Dimitri & Yorick 61 83.6% 88.5% 16 0
-Sokoban Jr. 1 60 80.0% 81.7% 2 19 73.7% 73.7% 0
-Howard’s 3rd set 40 60.0% 62.5% 4 1 0.0% 0.0% 0
-Simple sokoban 61 54.1% 60.7% 16 51 47.1% 54.9% 16
-Sokoban Jr. 2 54 48.1% 50.0% 2 40 45.0% 47.5% 2
-Microban 155 23.9% 24.5% 4 17 5.9% 11.8% 2
-Deluxe 55 21.8% 23.6% 2 1 0.0% 0.0% 0
-Sokogen 990602 78 20.5% 23.1% 8 0
-Yoshio Automatic 52 13.5% 17.3% 2 0
-Sasquatch III 16 6.2% 6.2% 0 8 0.0% 0.0% 0
-Howard’s 1st set 100 5.0% 6.0% 2 54 0.0% 0.0% 0
-Howard’s 2nd set 40 5.0% 10.0% 4 22 0.0% 0.0% 0
-Microcosmos 40 5.0% 7.5% 4 0
-Still more levels 35 2.9% 2.9% 0 34 2.9% 2.9% 0
-Sasquatch IV 36 2.8% 2.8% 0 20 0.0% 0.0% 0
-Xsokoban 40 2.5% 5.0% 128 39 2.6% 5.1% 128
-Sasquatch 49 2.0% 4.1% 4 39 0.0% 2.6% 4
+The “max solved” columns represent the proportion of levels solved at the number of steps in the “max at” column, which is the highest solved proportion for each number of thinking steps tried.
+The “largest level size” column represents the height and width of the solved level with the largest grid area.
+All levels Levels larger than $10\times 10$ Largest
+Level collection # solved Max slv. max at # solved Max slv. max at level size
+Dimitri & Yorick 61 86.9% 93.4% 2 0 (10, 12)
+Sokoban Jr. 1 60 85.0% 85.0% 0 19 73.7% 73.7% 0 (18, 21)
+Howard’s 3rd set 40 70.0% 70.0% 0 1 0.0% 0.0% 0 (11, 10)
+Simple sokoban 61 55.7% 62.3% 16 51 47.1% 54.9% 16 (16, 19)
+Sokoban Jr. 2 54 48.1% 53.7% 2 40 45.0% 47.5% 2 (14, 27)
+Sokogen 990602 78 37.2% 43.6% 4 0 (10, 10)
+Microban 155 31.0% 31.6% 64 17 5.9% 11.8% 2 (14, 10)
+Yoshio Automatic 52 28.8% 36.5% 2 0 (10, 10)
+Deluxe 55 25.5% 27.3% 16 1 0.0% 0.0% 0 (10, 13)
+Howard’s 2nd set 40 12.5% 15.0% 2 22 0.0% 0.0% 0 (11, 10)
+Sasquatch III 16 6.2% 6.2% 0 8 0.0% 0.0% 0 (10, 17)
+Microcosmos 40 5.0% 10.0% 16 0 (10, 10)
+Howard’s 1st set 100 4.0% 4.0% 0 54 0.0% 0.0% 0 (12, 10)
+Still more levels 35 2.9% 2.9% 0 34 2.9% 2.9% 0 (13, 13)
+Sasquatch IV 36 2.8% 2.8% 0 20 0.0% 0.0% 0 (10, 10)
+Xsokoban 40 2.5% 5.0% 128 39 2.6% 5.1% 128 (13, 19)
+Sasquatch 49 2.0% 4.1% 4 39 0.0% 2.6% 4 (14, 24)
 David Holland 1 10 0.0% 0.0% 0 5 0.0% 0.0% 0
@@
 Sokoban 50 0.0% 0.0% 0 48 0.0% 0.0% 0
-Appendix C Case Studies
+C.1 Generalization to zig-zag levels
+Figure 16: $16\times 16$ zig-zag level that the $\text{DRC}(3,3)$ doesn’t solve.
+The network solves all levels of this pattern only up to $15\times 15$ size and fails on all zig-zag levels beyond that.
+We create a template of levels as shown in Figure 16 where the agent has to take boxes from the left side to the targets on the right side by going through multiple zig-zag vertical lanes.
+This template can be created for arbitrary sized levels by adjusting the number of vertical lanes that the agent has to pass through.
+We find that the agent is able to solve levels only until the size of $15\times 15$ with 5 lanes but fails to solve larger levels.
+In larger levels, the agent does push the boxes through some of the lanes but is unable to find the complete path to the target through the remaining zig-zag lanes.
+Appendix D Bistable and unstable plans in toy environments
+Figure 17: Left: A custom level where the agent has two equally good paths to follow.
+The arrows show the prediction of the Agent-Directions probe with opacity proportional to the number of times an arrow was predicted across all the steps.
+The green and red arrows are correct and incorrect predictions, respectively.
+The agent behaves suboptimally by returning to the start and going left after first going right for three steps.
+Middle: The agent takes the path on the right after intervening with the corresponding arrows using the Agent-Directions probe on the first step.
+The same happens on the left if we intervene on that path without the suboptimal steps right of the undisturbed agent.
+Right: An empty level with Box-Directions probe intervening on the first four steps which are followed correctly by the agent on those four steps.
+When the intervention is removed on the fifth step, the agent computes the simpler path in green and doesn’t follow the path laid out earlier (in red).
+We examined the Box-Directions probe, which displayed significantly higher causal influence than the Agent-Directions probe Table 1.
+To better understand this behavior, we designed controlled scenarios.
+Figure 17 (left) shows a level where a single box is next to a target, and the agent faces two equally viable paths.
+Without intervention, the agent initially chooses the right, taking three steps before returning to the start and switching to the left.
+This behavior deviates from optimality, which requires committing to a single path.
+When we applied Agent-Directions probe to enforce a path choice, the agent consistently followed the chosen trajectory Figure 17 (middle).
+However, in levels where multiple viable paths exists, the Agent-Directions probe’s influence diminished, and the Box-Directions probe guided the agent more effectively.
+Figure 17 (right) shows a near-empty level where the Box-Directions probe was used to guide the agent on the first four steps.
+After the intervention ceased.
+the agent computed a shorter, optimal path and deviated from the initial intervention.
+This analysis the agent prioritizes box-directions when available and relies on the agent-direction cues as secondary guidance.
+These findings highlight the hierarchy of planning components encoded in the network.
+Appendix E Case Studies
 (a) file 0, level 18
@@
 Figure 18: Case studies of three medium-validation levels demonstrating different behaviors after 6 thinking steps.
-Colors are as in Figure 12.
-Boxes and targets are paired in upper- and lower-case letters respectively, such that the DRC’s best solution places boxes on targets in alphabetical order.
+Colors are as in Figure 1 (right).
+Boxes and targets are paired in upper- and lower-case letters respectively, and the optimal solution places boxes alphabetically.
 Videos available at this https URL.
-Levels solved faster obtain higher return as the agent incurs the per-step penalty fewer times.
-Note that the letters used are for reference and not a property of Sokoban.
+Levels solved faster incur fewer per-step penalties, yielding higher returns.
+Letters are for reference only and not intrinsic to Sokoban.
 Case Study: Thinking makes some levels solvable Figure 18(a).
-For example, thinking lets the DRC solve Figure 10(a).
-In the no-thinking condition, the DRC first pushes box $C$ one square to the right.
-It then goes back to push $A$ to $a$, but it has now become impossible to push box $B$ onto $b$.
-In contrast, after thinking, the DRC pushes $A$ to $a$ first, which lets it solve the level.
+In this scenario, the DRC fails to solve the level in the no-thinking condition.
+It initially pushes box $C$ one square to the right.
+While attempting to push $A$ to $a$, it blocks $B$ from reaching $b$, making the level unsolvable.
+In the thinking condition, the DRC correctly pushes $A$ to $a$ first, leaving room to push $B$ to $b$ and $C$ to ts target, solving the level.
 Case Study: Thinking speeds up solving Figure 18(b).
-Figure 10(b) illustrates a scenario where the agent happens.
-In the no-thinking condition, the DRC takes many steps back and forth before pushing any boxes.
-First it goes down to $y$, up to $c$, then down onto $z$, back up to $y$ and to $z$ again.
-It then proceeds to solve the rest of the puzzle: push box $A$ onto $a$, prepare box $B$ on $x$ and box $C$ where $B$ originally was, push in boxes $B,C$ and finally $D$.
-In the thinking condition, the DRC makes a beeline for $A$ and then plays the same solution.
+Without thinking, the DRC takes inefficient route, moving back and forth multiple times before starting to push boxes.
+For example, it moves to $y$, up to $c$, then down onto $z$, back up to $y$ and to $z$ before solving the puzzle.
+With thinking, the DRC immediately moves to box $A$, completing the solution with fewer extraneous steps.
+The resulting solution is faster, reducing the total step penalty.
 Case Study: Thinking slows down Figure 18(c).
-Figure 10(c) illustrates a scenario where thinking time results in a slower solution.
-In the no thinking condition, the DRC starts by pushing box C into position y, then pushes boxes A, then B into place.
+In this level, thinking causes the DRC to take a less efficient path.
+Without thinking, the DRC pushes box $C$ to $y$, followed by boxes $A$, then $B$ into place.
 On the way back down, the DRC pushes $C$ onto $c$ and finally $D$ onto $d$.
-In contrast, in the thinking condition, the DRC goes the other way and starts by pushing B onto b.
-The subsequent solution (A, C, then D) is the same, but the DRC has wasted time trekking back from B to A.
-Appendix D Probe Training
+In the thinking condition, the DRC starts by pushing B onto b then backtracks to push A, C, then D.
+This unnecessary backtracking leads to a slower overall solution.
+Appendix F Probe Training
 Probe architecture
-We train multiple different linear probes on the hidden states $h$ and cell states $c$ activations of the network.
-The probes are trained either on the states of the individual layers or by concatenating the states of all the three layers.
-We collect the activations of the model by letting it play through the hard levels.
-The training set for the probes are constructed by taking a random sample of cached levels and including all the timesteps in the level excluding the first five timesteps.
-We exclude these intial steps as we believe they can be noisy as the network comes up with the stategy to solve the levels.
-The probes are trained with logistic regression with L1 decay using the Scikit-Learn library.
-We do a grid-search on the learning rate and the L1 weight decay to select the probe that has the highest F1-score on a validation set.
-For multi-class probe targets, each potential output class $l$ is treated like a different data point.
-That is: for each label $l$, we consider a prediction positive if the highest probe logit is for label $l$, and negative otherwise.
-We consider a data point positive iff the true label is label $l$.
-Then, we compute the confusion matrix and the F1 score from the $n\cdot l$ data points.
-Appendix E Details about Sparse Autoencoders (SAEs)
+We train multiple linear probes on the hidden states $h$ and cell states $c$ activations of the network.
+Probes are trained either on activations from individual layers or by concatenating activations across all three layers.
+Activations were collected as the model played through hard levels.
+The training set are constructed by randomly sampling cached levels and including all timesteps except the first five.
+These initial steps excluded due to potential noise while the network formulated its strategy for solving the levels.
+Probes are trained with logistic regression with L1 decay using the Scikit-Learn library.
+A grid search over learning rates and L1 weight decay was conducted to identify the probe with the highest F1 score on the validation set.
+For multi-class probe targets, each potential output class $l$ is treated as a separate data point.
+Specificially: A prediction was considered positive if the highest probe logit corresponded to class $l$, and negative otherwise.
+A data point was positive if its true label matched $l$.
+The confusion matrix and F1 score were then computed from the $n\cdot l$ data points.
+In addition to Agent-Directions and Box-Directions probes, we trained the following:
+  • Next-Box probe: Predicts $1$ on the square of the box that the agent will move next and 00 for every other square.
+  • Next-Target probe: Predicts $1$ on the square of the target that the agent will put a box in next, and 00 for every other square.
+  • Next-Action probe: Grid-wise binary probe for each of the four move action that predict $1$ for all squares in the grid if the next action matches the probe’s action.
+We train grid-wise probe instead of a global probe so that the probe can be applied to arbitrary-sized inputs.
+The results for the next-action probes are available in Table 10.
+  • Pacing probe: The global label is 1 if the agent is currently in a cycle, and 0 otherwise.
+  • Value probe: The global label is the numerical value that the critic head outputs.
+Probe results.
+The pacing probe gets $F_{1}=31.0\%$, which is not much better than the constant 1 probe, which has $F_{1}=12.8\%$.
+This suggests that the $\text{DRC}(3,3)$ does not represent whether it is in a cycle or not.
+For the value function probe, we compute the fraction of variance explained $R^{2}$.
+If we train a global probe, $R^{2}$ is very high: $97.7\%$$99.7\%$ depending on the layer.
+However, grid-wise probes obtain much worse but still passable results: $41.0\%$$79.2\%$ depending on layer.
+We visually checked whether the grid-wise probe reads off the values of different plans, but could not find any such pattern.
+Almost all the performance of the global probe is recovered by training on the mean-pooled inputs: $95.2\%$$99.5\%$.
+It is likely that the global and mean-pooled value probes are indirectly counting the number of squares the agent will step on, which almost fully determines the value, and we know is possible due to the future-direction probes.
+The possible box directions probe got approximately $\approx 30\%$ accuracy, and it was
+Table 5: Weight and bias for transforming action probes to predictions
+Mean Max Positive proportion
+Weight 1.2086 -0.0582 0.2070
+Up Down Left Right
+Bias 0.3337 -0.0921 -0.0632 -0.0539
+Appendix G Looking for interpretable features with Sparse Autoencoders (SAEs)
 Figure 19: Visualization of some interpretable features from the SAE of last layer.
@@
 The precision, recall, and F1 score for the features are reported in Table 9.
-In order to search for more monosemantic and interpretable features in the network, we train sparse-autoencoders (SAEs) [Huben et al., 2023, Bricken et al., 2023] on the individual squares in the $h$ hidden state of the network consisting of 32 neurons.
+To search for monosemantic and interpretable features in the network, we train sparse-autoencoders (SAEs) [Huben et al., 2023, Bricken et al., 2023] on the individual squares in the $h$ hidden state of the network consisting of 32 neurons.
 Thus, we get a $10\times 10$ visualization for each SAE feature as shown in Figure 19.
-We use the top-k activation function [Gao et al., 2024, Tamkin et al., 2024] for the SAE as it is currently the state-of-the-art method to train SAEs and is easier to work with as it directly sets the $L_{0}$ norm of the SAE activations.
-The hyperparameter search space for training the SAE is provided in Table 7.
-We train separate SAEs for each layer with the specified hyperparameters and pick the one that achieves greater than $90\%$ explained variance while having interpretable features assessed through manual visual inspection.
-We release these trained SAEs and probes in the same huggingface repo as our trained DRC networks.
+We use the top-k activation function [Gao et al., 2024, Tamkin et al., 2024], which is state-of-the-art for training SAEs and directly enforces an $L_{0}$ sparcity constraint on activations.
+The SAE hyperparameter search space is detailed in Table 6.
+We train separate SAEs for each layer with the specified hyperparameters, selecting those achieving greater than $90\%$ explained variance while retaining interpretable features based on manual visual inspection.
+We release these trained SAEs and probes in the huggingface repository with our trained DRC networks.
 [^8] https://huggingface.co/AlignmentResearch/learned-planner
-We provide examples of some of the interpretable features from an SAE trained on the last layer with $k=8$ in Table 8 with corresponding visualization in Figure 11.
-For the “Target”, “Unsolved”, and “Solved” concepts (Table 10), we observed cases where the SAE feature is offset from the ground truth by 1 square in either horizontal or vertical directions.
-The level setup includes a permanent one-square outer-edge wall, so this offset never results in an out-of-bounds issue.
+Examples of interpretable features from an SAE trained on the last layer with $k=8$ are provided in Table 7, with visualizations in Figure 19.
+For “Target”, “Unsolved”, and “Solved” concepts (Table 9), we occasionally observed an offset of 1 square (horizontal or vertical) from the ground truth.
+Due to a permanent one-square outer-edge wall, this offset never results in out-of-bounds errors.
 We evaluated these potential “Offset” variants for the Target, Unsolved, and Solved concepts.
-All of the features in the SAE that we find to be interpretable are already embedded in individual channels.
-These channels are either as monosemantic as the SAE features, or in some cases more monosemantic as measured by the F1 score.
-Table 3 reports the precision, recall, and F1 scores for the action features as evaluated through channels, SAE neurons, and linear probe trained against the ground truth action predictions.
-Table 10 reports the same scores for additional interpretable features.
-The SAE action features have lower F1 scores than channels by a margin of $5.9\%$ on average.
-On the other hand, the linear probes trained across all channels of the hidden state have similar F1 scores as the specified channels indicating that the channels are monosemantic and cannot be improved upon by combining various channels using a linear probe.
+All interpretable SAE features identified are already embedded in individual channels, with comparable or higher monosemanticity then the SAE features based on F1 scores.
+Table 10 compares precision, recall, and F1 scores for action features across channels, SAE neurons, and linear probes trained against the ground truth predictions.
+Table 9 reports scores for other interpretable features.
+SAE action features consistently underperform channels by an average F1 margin of $5.9\%$.
+Linear probes trained across all hidden-state channels achieve similar F1 scores to individual channels indicating the inherent monosemanticity of channels without further improvement through linear probe.
 Table 6: Hyperparameter search space for training SAE
@@
 Solved targets (0, 0) -L3C7 91.6 94.6 93.0 L3F179 91.7 91.5 91.6
-Appendix F Additional quantitative behavior figures and tables
-Figure 12: Left: A Sokoban level from the hard set. Right: the same level as the NN sees it, one pixel per tile. Walls are black, boxes are brown, targets are pink and the robot is green.
+Table 10: Action features scores across channels, probes, and SAE features
+Feature Channel SAE Feature Probe
+Number Prec Rec F1 Number Prec Rec F1 Prec Rec F1
+Up L3C29 95.7 88.1 91.7 L3F270 93.9 76.2 84.1 97.5 86.5 91.7
+Down L3C8 98.4 80.8 88.8 L3F187 98.0 79.1 87.6 97.6 86.9 91.9
+Left L3C27 85.5 84.6 85.1 L3F244 96.1 63.2 76.2 83.5 86.6 85.0
+Right L3C3 97.0 86.9 91.7 L3F385 94.6 78.5 85.8 97.6 87.4 92.2
+Appendix H Additional quantitative behavior figures and tables
+Figure 21 through Figure 23 present quantitative behavior, analyzing various aspects of the network’s planning behavior and evaluation metrics.
 Figure 20: Number of thinking steps required to solve the level vs.
 number of nodes A* needs to expand to solve it.
-The trend is increasing at the end but very unclear, indicating different heuristics used by the DRC and A*.
-Figure 14: Success rate on datasets of various difficulty, for various checkpoints of each architecture.
-This deviates very little from a curve, which shows that ResNets and DRCs which are equally good at the easier sets are also equally good at the harder sets.
-Perhaps DRC(1,1) is a slight exception, but it also performs much worse than the others overall (see Figure 7).
+The weak correlation toward the end indicates different that the DRC and A* rely on different heuristics.
+Figure 21: Success rate on datasets of varying difficulty for different architecture checkpoints.
+Performance trends suggest ResNets and DRCs with comparable results on easier sets also perform similarly on harder sets.
+DRC(1,1) is a slight exception, performing consistently worse overall (see Figure 8).
 (a) Cycle length distribution
 (b) with cycles from all levels
-(c) with cycles from those levels that were solved by the agent
-Figure 15: We replace $N$-length cycles with $N$ thinking steps and check for the same state after some timesteps.
-(a) A histogram of cycle lengths in the medium-validation set.
+(c) with cycles from levels solved by agent
+Figure 22: We replace $N$-length cycles with $N$ thinking steps to examine state consistncy across subsequent timesteps.
+(a) Histogram of cycle lengths in the medium-validation set.
 (b, c) After replacing a cycle with the same length in thinking steps, are all the states the same for the next $x$ steps?
+(a) Pacing behavior on file 0, level 53.
+On the given starting observation, the agent paces around 4 spaces in the first 9 steps and then goes on to solve the level.
+Video for the level is available at this https url .
+(b) Change in per-step F1 score of Box-Directions probe for moves in cycles and outside cycles on medium-difficulty validation levels.
+The non-cycle moves were recorded from the same distribution of timesteps where cycles occur but from levels without a cycle at those steps.
+Mean per-step change in F1 for cycle and non-cycle steps are $1.40\%\pm 0.06\%$ and $0.84\%\pm 0.04\%$ respectively.
+Figure 23: Illustration of cycles and F1 scores
+Appendix I Additional related work
+Ethical treatment of AIs.
+The question whether AIs deserve moral consideration has been widely debated.
+Schwitzgebel and Garza [2015] argue that highly human-like AIs merit rights, while Tomasik [2015] suggest that most AIs deserve some degree of consideration, akin to biological organisms [Singer, 2004].
+The concept of ethical treatement for reinforcement learners has been explored by Daswani and Leike [2015], who propose that pleasure and pain in these systems may correspond to temporal difference (TD) error rather than absolute returns.
+If neural networks have internal objectives distinct from their critic head [Hubinger et al., 2019, Di Langosco et al., 2022], identifying these objectives could provide higher-assurance methods for assessing AI goals compared to direct querying [Perez and Long, 2023].
+Chain-of-thought faithfulness.
+The faithfulness of chain of thought reasoning in large language models (LLMs).
+Studies such as [Lanham et al., 2023, Pfau et al., 2024] investigate whether LLMs rely on plain English for long-term reasoning, which could allow unintended consequences to be easily identified and mitigated Scheurer et al.
+[2023].
+Fully reverse engineering small networks.
+Recent efforts have successfully reverse-engineered small neural networks performing algorithmic tasks [Nanda et al., 2023a, Chughtai et al., 2023, Zhong et al., 2023, Quirke and Barez, 2023].
+Systematic Generalization.
+Previous work identified conditions under which neural networks generalize, such as diverse datapoints and egocentric environments [Lake and Baroni, 2023, Hill et al., 2020, Mutti et al., 2022].
+Similar interpretability can be extended across these neural networks to uncover shared planning mechanisms and the conditions in which they emerge.
+Appendix J Sussman’s anomaly
+Although the $\text{DRC}(3,3)$ shows exceptional long-term planning capabilities as demonstrated in the paper, it can have some trivial failure modes.
+Figure 24 shows a level in which the agent indefinitely tries to put the two boxes on the right onto the same target that is on the right side of the level.
+This failure mode is similar to the Sussman’s anomaly demonstrated by Sussman [1973] illustrating the weakness of non-interleaved planning algorithms.
+We observe that such cases happen rarely, with the network being able to resolve the interleaved dependencies between boxes in most cases.
+Figure 24: Level 466 from file 233 of train-medium in which the $\text{DRC}(3,3)$ network tries to put both the boxes on the right to the target on the right by repeatedly putting one box on target, only to remove it and put the other one on the same target.
+Appendix K Planning vs. predicting box-directions
+Most of our analysis and experimentation in the paper relies on the box-directions probe, which predicts the sequence of all future box movement directions from squares in the grid given network’s activations.
+While predicting the box-directions is very close to planning, it misses a few details required for a complete plan.
+For example, for a complete plan of actions, the agent also has to also know the order in which it wants to move the boxes in and the plan a path from the agent’s current position to the next box to push.
+However, we do train probes for agent-directions and next-box to move and show they have high predictive accuracy.
+We focus primarily on boxes-directions because it is the most crucial component of a plan in the game, which is also reflected in the fact that the box-directions probe is the most causal probe.